Hallc beamtest runlist » History » Version 6
Jixie Zhang, 10/20/2022 03:54 PM
| 1 | 2 | Jixie Zhang | #This file record the run list for solid ecal beamtest in HallC. |
|---|---|---|---|
| 2 | 1 | Jixie Zhang | #Do not delete |
| 3 | |||
| 4 | 5 | Jixie Zhang | 5-pass beam, 10.549 GeV |
| 5 | |||
| 6 | ================================================================= |
||
| 7 | 2 | Jixie Zhang | slot 7 FADC250_DAC 3300 (for all 16 channels) |
| 8 | slot 8 FADC250_DAC 3300 (for all 16 channels) |
||
| 9 | |||
| 10 | 20220620 21:30 |
||
| 11 | HV slot 2: 0 0 0 0 0 0 0 |
||
| 12 | HV slot 10: 1500 1500 1500 1500 0 0 0 0 0 0 0 |
||
| 13 | |||
| 14 | 2925 FADC only |
||
| 15 | 2926 FADC only |
||
| 16 | 2927 FADC only |
||
| 17 | 2928 FADC only |
||
| 18 | 2929 FADC only |
||
| 19 | |||
| 20 | File not exist for run 2930-2949, debug DAQ |
||
| 21 | |||
| 22 | 20220621 12:20 |
||
| 23 | HV slot 2: 900 950 900 0 0 0 0 |
||
| 24 | HV slot 10: 1500 1500 1500 1500 1500 1500 1500 1800 1800 1800 |
||
| 25 | |||
| 26 | 2950-2956 SSP only |
||
| 27 | 2957 SSP only, trash |
||
| 28 | 2958 debug DAQ, trash |
||
| 29 | 2959 First run for FADC+SSP, good data |
||
| 30 | 2960 good FADC+SSP data |
||
| 31 | 2961-2964 DEBUG SSP, all trash |
||
| 32 | |||
| 33 | |||
| 34 | 20220623 15:50 |
||
| 35 | HV slot 2: 900 950 900 1500 1500 1500 1500 |
||
| 36 | HV slot 10: 1600 1600 1700 1600 1500 1500 1500 1800 1800 1800 |
||
| 37 | |||
| 38 | From this run, trigger lost due to FIFO has an 100mv offset |
||
| 39 | 2965-2968 SSP only, no trigger and debug daq, trash |
||
| 40 | 2969 FADC+SSP, no trigger, trash |
||
| 41 | 2970 cosmic run, no trigger, trash |
||
| 42 | 2971 beam run, no trigger, trash |
||
| 43 | |||
| 44 | From run 2972, trigger is connected to another FIFO unit which has 0 offset |
||
| 45 | 20220624 12:20 |
||
| 46 | HV slot 2: 900 950 900 1000 1000 1000 1000 |
||
| 47 | HV slot 10: 1600 1600 1700 1600 1500 1500 1500 1800 1800 1800 |
||
| 48 | 2972 debug DAQ, trash |
||
| 49 | 2973 good FADC+SSP data <------- First run for good Cherenkov HV |
||
| 50 | 2974 |
||
| 51 | 2975 trash |
||
| 52 | 2976-2981 good FADC+SSP data |
||
| 53 | 2982 trash |
||
| 54 | 2983-3002 good FADC+SSP data |
||
| 55 | 3003-3005 debug DAQ, trash |
||
| 56 | 3006-3013 good FADC+SSP data |
||
| 57 | 3014-3015 cosmic |
||
| 58 | 3016 daq crash, no output files |
||
| 59 | 3017-3018 debug daq, trash |
||
| 60 | 3 | Jixie Zhang | 3019 cosmic |
| 61 | 4 | Jixie Zhang | 3020-3021 debug daq for DAC values and HV, trash |
| 62 | ===================================================================== |
||
| 63 | 2 | Jixie Zhang | GOOD DATA STARTED ...... |
| 64 | 3 | Jixie Zhang | ===================================================================== |
| 65 | 2 | Jixie Zhang | For all runs before 3022, slot 7 and slot 8 FADC250_DAC are 3300 for all channels. Therefore some channels might not see signal. |
| 66 | |||
| 67 | 20220629 16:20 |
||
| 68 | HV slot 2: 900 950 900 1000 1000 1000 1000 |
||
| 69 | HV slot 10: 1600 1600 1850 1750 1000 1100 1800 1680 1920 1800 |
||
| 70 | |||
| 71 | slot 8 FADC250_DAC 3000 |
||
| 72 | slot 8 FADC250_DAC 3300 3000 3300 3000 3000 3000 3000 3000 3000 3000 3000 3000 3000 3000 3000 3000 |
||
| 73 | |||
| 74 | 3022 First production run <----------------- |
||
| 75 | 3023 DAQ crashed |
||
| 76 | 3024 production |
||
| 77 | 3025 debug run, cosmic, trash |
||
| 78 | 4 | Jixie Zhang | |
| 79 | 2 | Jixie Zhang | ===================================================================== |
| 80 | 4 | Jixie Zhang | Cherenkov ch12 (started from 0) has cabling issue for runs before 3026 |
| 81 | 2 | Jixie Zhang | ===================================================================== |
| 82 | |||
| 83 | Change FADC DAC for Ecal channels |
||
| 84 | slot 8 FADC250_DAC 3300 3000 3300 3000 3000 3000 3000 3000 3000 3000 3300 3300 3300 3000 3300 3000 |
||
| 85 | |||
| 86 | 3026 production, cosmic at the first 4 hours(~30 events), see beam around 6 pm |
||
| 87 | 3027 production |
||
| 88 | 3028 production |
||
| 89 | 3029 production, no beam at the beginning (several hours) |
||
| 90 | 3030 production |
||
| 91 | 3031 production |
||
| 92 | 3032 production |
||
| 93 | 3033 production |
||
| 94 | 3034 production |
||
| 95 | 3035 production |
||
| 96 | 3036 production |
||
| 97 | 3037 production, daq machine tunl11 has some issue and vncserver crashed, so coda also crashed |
||
| 98 | 3038 production, first test run after rebooting tunl11 and sbsvme22, data is good |
||
| 99 | 3039 production, test auto log scripts, data is good |
||
| 100 | 3040 production |
||
| 101 | 3041 cosmic, trash |
||
| 102 | 3042 debug daq,trash |
||
| 103 | 3043 no files, daq crashed |
||
| 104 | |||
| 105 | |||
| 106 | FADC config file changed: |
||
| 107 | 1) remove slot 9 |
||
| 108 | 2) slot 7 |
||
| 109 | |||
| 110 | FADC250_ALLCH_DAC 3205 3170 3170 3195 3195 3190 3200 3195 3180 3205 3190 3215 3170 3215 3175 3205 |
||
| 111 | |||
| 112 | 3)slot 8 |
||
| 113 | |||
| 114 | FADC250_W_WIDTH 400 |
||
| 115 | |||
| 116 | FADC250_ALLCH_DAC 3500 3200 3460 3190 3220 3210 3190 3220 3220 3190 3500 3520 3510 3500 3490 3450 |
||
| 117 | |||
| 118 | |||
| 119 | 3044 cosmic, slot 8 FADC250_W_WIDTH 400 |
||
| 120 | 3045 cosmic, slot 8 FADC250_W_WIDTH 400 |
||
| 121 | |||
| 122 | |||
| 123 | 1 | Jixie Zhang | changed slot 7 FADC250_W_WIDTH from 200 to 400 |
| 124 | |||
| 125 | 3046 both slot 7 and slot 8, FADC250_W_WIDTH = 400 |
||
| 126 | 3047 production |
||
| 127 | 2 | Jixie Zhang | 3048 production |
| 128 | 3049 1 event only, cosmic, end it for changing config file |
||
| 129 | |||
| 130 | ============================================================== |
||
| 131 | 5 | Jixie Zhang | :::::::::::::::::::::::USEFUL DATA START:::::::::::::::::::::: |
| 132 | Before run 3050, only part of the pulse was recorded, therefore all these runs are gabages. |
||
| 133 | Runs 3050 - 3057 could be useful, but their HVs are not perfect. |
||
| 134 | ============================================================== |
||
| 135 | for both slot #7 and #8, changed to the following 2 lines: |
||
| 136 | 2 | Jixie Zhang | FADC250_W_OFFSET 1000 |
| 137 | FADC250_W_WIDTH 1000 |
||
| 138 | |||
| 139 | 3050 debug fadc config file,FADC250_W_OFFSET=1000 and FADC250_W_WIDTH=1000 |
||
| 140 | 3051 debug fadc config file,FADC250_W_OFFSET=1000 and FADC250_W_WIDTH=1000 |
||
| 141 | 1 | Jixie Zhang | |
| 142 | 3 | Jixie Zhang | ============================================================== |
| 143 | 2 | Jixie Zhang | for both slot #7 and #8, changed to the following 2 lines: |
| 144 | FADC250_W_OFFSET 1060 |
||
| 145 | FADC250_W_WIDTH 400 |
||
| 146 | |||
| 147 | 3052 debug fadc config file,FADC250_W_OFFSET=1060 and FADC250_W_WIDTH=400 |
||
| 148 | 3053 production |
||
| 149 | 3054 production |
||
| 150 | |||
| 151 | 4 | Jixie Zhang | ============================================================== |
| 152 | 1 | Jixie Zhang | HV changed |
| 153 | Slot 2 |
||
| 154 | 2 | Jixie Zhang | Name: Shower_L(SDU#4), Shower_R(THU#2), Shower_T(SDU#5), CherA, CherB, CherC, CherD |
| 155 | HV: 950 980 900 1000 1000 1000 1000 |
||
| 156 | |||
| 157 | Slot 10 |
||
| 158 | Name: SC0_T, SC0_B, SC1_T, SC1_B, LASPD_T, LASPD_B, SPD, Preshower_L(SDU#1), Preshower_R(THU#1), Preshower_T(SDU#2) |
||
| 159 | HV: 1600 1600 1750 1750 900 1000 1950 1580 1750 1750 |
||
| 160 | |||
| 161 | 1 | Jixie Zhang | 3055 HV tripped, empty,trash |
| 162 | 3056 production |
||
| 163 | 3057 production |
||
| 164 | |||
| 165 | ================================================================ |
||
| 166 | 5 | Jixie Zhang | ::::::::::::::::::::::::GOODEN DATA START::::::::::::::::::::::: |
| 167 | ================================================================ |
||
| 168 | 1 | Jixie Zhang | HV changed |
| 169 | Slot 2 |
||
| 170 | Name: Shower_L(SDU#4), Shower_R(THU#2), Shower_T(SDU#5), CherA, CherB, CherC, CherD |
||
| 171 | HV: 1000 1030 900 1000 1000 1000 1000 |
||
| 172 | |||
| 173 | Slot 10 |
||
| 174 | Name: SC0_T, SC0_B, SC1_T, SC1_B, LASPD_T, LASPD_B, SPD, Preshower_L(SDU#1), Preshower_R(THU#1), Preshower_T(SDU#2) |
||
| 175 | HV: 1600 1600 1750 1750 880 1000 1950 1550 1800 1750 |
||
| 176 | 5 | Jixie Zhang | 3058 production, 05:15 July 9th - 03:50, target=Loop 2 LH2,shms_p=8.035,shms_theta=7.83,hms_p=-2.398,hms_theta=28.19, |
| 177 | 13:10 target changed to dummy, I=38uA, |
||
| 178 | 18:43 target changed to Loop 2 LH2, I=38uA |
||
| 179 | |||
| 180 | 3059 production, 03:56-15:24 July 10th,target=Loop 2 LH2,shms_p=8.035,shms_theta=7.83,hms_p=-2.398,hms_theta=28.19, |
||
| 181 | 3060 production, 15:25-00:35, target=Loop 2 LH2,shms_p=8.035,shms_theta=7.83,hms_p=-2.398,hms_theta=28.19, |
||
| 182 | 3061 production, 00:36-09:41 July 11th,target=Loop 2 LH2,shms_p=8.035,shms_theta=7.83,hms_p=-2.398,hms_theta=28.19, |
||
| 183 | 05:55 SHMS rotated to 9.83deg, shms_p=8.035,shms_theta=9.83,hms_p=-2.398,hms_theta=28.19, |
||
| 184 | |||
| 185 | 3062 production, 09:41-10:15 July 11th,target=Loop 2 LH2,shms_p=8.035,shms_theta=9.83,hms_p=-2.398,hms_theta=28.19, |
||
| 186 | 1 | Jixie Zhang | 3063 cosmic, 6 events |
| 187 | 5 | Jixie Zhang | |
| 188 | 12:53 SHMS rotated to 5.65deg |
||
| 189 | |||
| 190 | 3064 production, 16:06-11:19 July 12th,target=Loop 2 LH2,shms_p=8.035,shms_theta=5.65,hms_p=-2.398,hms_theta=28.19 |
||
| 191 | |||
| 192 | 3065 production, 11:19-18:02 July 12th,target=Loop 2 LH2,shms_p=8.035,shms_theta=5.65,hms_p=-2.398,hms_theta=28.19 |
||
| 193 | 3066 production, 18:02-00:43,target=Loop 2 LH2,shms_p=8.035,shms_theta=5.65,hms_p=-2.398,hms_theta=28.19 |
||
| 194 | 3067 production, 00:45-11:48 July 13th,target=Loop 2 LH2,shms_p=8.035,shms_theta=5.65,hms_p=-2.398,hms_theta=28.19 |
||
| 195 | 3068 production, 11:48-02:04, target=Loop 2 LH2,shms_p=8.035,shms_theta=5.65,hms_p=-2.398,hms_theta=28.19 |
||
| 196 | 13:15, target ladder moved to home, no target (absolute empty) |
||
| 197 | 14:50, target changed to dummy |
||
| 198 | 23:28, target changed to Loop 2 10cm LH2 |
||
| 199 | |||
| 200 | 3069 production, 02:04-09:54 July 14th, target=Loop 2 LH2,shms_p=8.035,shms_theta=5.65,hms_p=-2.398,hms_theta=28.19 |
||
| 201 | 3070 production, 09:55 July 14th - 10:16 July 15th, target=Loop 2 LH2,shms_p=8.035,shms_theta=5.65,hms_p=-2.398,hms_theta=28.19 |
||
| 202 | 3071 production, 10:16-03:26 July 15th, target=Loop 2 LH2,shms_p=8.035,shms_theta=5.65,hms_p=-2.398,hms_theta=28.19 |
||
| 203 | 16:38 target changed to dummy |
||
| 204 | 19:33 target changed to Loop 2 LH2 |
||
| 205 | |||
| 206 | 3072 production, 03:27-16:39 July 16th, target=Loop 2 LH2,shms_p=8.035,shms_theta=5.65,hms_p=-2.398,hms_theta=28.19 |
||
| 207 | 13:29 SHMS rotated to 9.65deg, tried to get I=80uA |
||
| 208 | |||
| 209 | 3073 production, 16:42-14:26 July 17th,target=Loop 2 LH2,shms_p=8.035,shms_theta=9.66,hms_p=-2.398,hms_theta=28.19 |
||
| 210 | 3074 production, 14:26-02:11,target=Loop 2 LH2,shms_p=8.035,shms_theta=9.66,hms_p=-2.398,hms_theta=28.19 |
||
| 211 | 3075 production, 02:11-10:35 July 18th,target=Loop 2 LH2,shms_p=8.035,shms_theta=9.66,hms_p=-2.398,hms_theta=28.19 |
||
| 212 | 3076 production, 10:16-00:39, target=Loop 2 LH2,shms_p=8.035,shms_theta=9.66,hms_p=-2.398,hms_theta=28.19 |
||
| 213 | 3077 production, 00:39-09:42 July 19th, target=Loop 2 LH2,shms_p=8.035,shms_theta=9.66,hms_p=-2.398,hms_theta=28.19 |
||
| 214 | 06:51 target changed to dummy |
||
| 215 | |||
| 216 | 3078 production, 09:43-00:29 July 20th, target=dummy,shms_p=8.035,shms_theta=9.66,hms_p=-2.398,hms_theta=28.19 |
||
| 217 | 10:04 target changed to Loop 2 LH2, target=Loop 2 LH2,shms_p=8.035,shms_theta=9.66,hms_p=-2.398,hms_theta=28.19 |
||
| 218 | 12:00-ish, SHMS rotated to 8.68deg, HMS rotated to 34.15deg, |
||
| 219 | target_type=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 220 | |||
| 221 | beam stopped at 19:00-ish, no beam for whole day of July 20th |
||
| 222 | =========================================================== |
||
| 223 | 1 | Jixie Zhang | 3079 pulse trigger |
| 224 | 3080 pulse trigger |
||
| 225 | 3081 pulse trigger |
||
| 226 | 3082 pulse trigger |
||
| 227 | 5 | Jixie Zhang | =========================================================== |
| 228 | 1 | Jixie Zhang | 3083-3084 trash, no data file |
| 229 | 3085 debug daq |
||
| 230 | 5 | Jixie Zhang | 3086 production, 03:11-09:42 July 21st, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
| 231 | 3087 production, 09:43-16:07 July 21st, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 232 | 3088 production, 16:07-12:23 July 22nd, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 233 | 3089 production, 12:23-04:28 July 23rd, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 234 | 3090 production, 04:28-15:44 July 23rd, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 235 | 3091 production, 15:47-05:28 July 24th, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 236 | 3092 production, 05:28-18:22 July 24th, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 237 | 3093 production, 18:22-09:32 July 25th, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 238 | 3094 production, 09:33-23:40 July 25th, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 239 | 3095 production, 23:41-09:29 July 26th, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 240 | 3096 production, 09:29-15:35,July 26th, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 241 | 13:45 target changed to dummy |
||
| 242 | beam stopped at ~15:10 due to thunderstorm. |
||
| 243 | |||
| 244 | 3097 production, 15:35 July 26th - 04:56 July 28th, target=dummy,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 245 | Dummy target when this run was started. At 22:45 July 26, target ladder was changed to home(empty). |
||
| 246 | Tune beam cane at 11:27 July 27th. |
||
| 247 | 12:15 July 27th, beam was stable to reach 65uA, changed target to Loop 2 LH2. |
||
| 248 | target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15 |
||
| 249 | 17:06 July 27th, changed target to dummy. |
||
| 250 | 22:06 July 27th, changed target to Loop 2 LH2. |
||
| 251 | |||
| 252 | 3098 production, 04:56-09:48 July 28th, target=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15 |
||
| 253 | |||
| 254 | updating readout list ...... |
||
| 255 | Bryan added a feature to run the fixed pulser at some frequency until some number of sync events. |
||
| 256 | It's currently configured (ti_list.c) with the structure: |
||
| 257 | |||
| 258 | typedef struct |
||
| 259 | { |
||
| 260 | int enable; // enable (1) / disable (0) |
||
| 261 | int interval; // 1 sync event every *interval* events |
||
| 262 | int pedestal; // disable (0) / Disable after *pedestal* sync events |
||
| 263 | int current; // sync event counter for pedestal feature |
||
| 264 | int pulser_arg1; // fixed pulser argument. Units of 120ns. |
||
| 265 | } TI_SYNCEVENT_CONFIG; |
||
| 266 | |||
| 267 | TI_SYNCEVENT_CONFIG tiSyncEventConfig = {1, 1000, 100, 0, 83}; |
||
| 268 | |||
| 269 | using hallc_ti configuation, with new ti_list.c |
||
| 270 | HV is off |
||
| 271 | 3100 pulse, HV off, beam off |
||
| 272 | 3101 pulse, HV off, beam off |
||
| 273 | 3102 pulse, HV off, beam off |
||
| 274 | |||
| 275 | HV is on |
||
| 276 | 3103 pulse, HV on, beam on |
||
| 277 | 3104 pulse, HV on, beam on |
||
| 278 | 3105 pulse, HV on, beam on |
||
| 279 | |||
| 280 | ========================================================================== |
||
| 281 | go back to use hallc_fadc_ssp configuration, with new ti_list.c |
||
| 282 | 3106 production, 11:41-12:19, target_type=Loop 2 10cm LH2,shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 283 | |||
| 284 | FADC config changed to make pedestal arround 100 |
||
| 285 | slot 7: |
||
| 286 | #FADC250_ALLCH_DAC 3205 3170 3170 3195 3195 3190 3200 3195 3180 3205 3190 3215 3170 3215 3175 3205 |
||
| 287 | FADC250_ALLCH_DAC 3150 3120 3120 3150 3150 3140 3150 3150 3130 3160 3140 3270 3120 3170 3130 3160 |
||
| 288 | |||
| 289 | slot 8: |
||
| 290 | #FADC250_ALLCH_DAC 3500 3200 3460 3190 3220 3210 3190 3220 3220 3190 3500 3520 3510 3500 3490 3450 |
||
| 291 | FADC250_ALLCH_DAC 3450 3150 3390 3140 3160 3150 3140 3170 3170 3130 3450 3470 3460 3450 3440 3390 |
||
| 292 | |||
| 293 | 3107-3111 debug daq, could be good for pedestal |
||
| 294 | 3112 production, daq crashed, could be good for pedestal |
||
| 295 | |||
| 296 | 3113 no data file, coda can not start |
||
| 297 | |||
| 298 | ========================================================================== |
||
| 299 | Reboot tunl11 and sbsvme22 |
||
| 300 | |||
| 301 | 3114 production, 03:02-16:52,target_type=dummy, shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 302 | 11:21, target changed to Loop 2 10cm LH2, shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 303 | |||
| 304 | 3115 production, 16:53-14:10 July 31st,target=Loop 2 10cm LH2, shms_p=7.913,shms_theta=8.68,hms_p=-2.341,hms_theta=34.15, |
||
| 305 | 17:14 SHMS rotated to 10.86deg, HMS rotated to 13.43deg, |
||
| 306 | beam came back at 19:35, target=Loop 2 10cm LH2, shms_p=6.265,shms_theta=10.86,hms_p=-4.256,hms_theta=13.43, |
||
| 307 | 01:01 target changed to dummy, |
||
| 308 | 01:50 SHMS rotated to 8.86 deg, |
||
| 309 | 01:53 target changed back to Loop 2 10cm LH2, shms_p=6.265,shms_theta=8.86,hms_p=-4.256,hms_theta=13.43, |
||
| 310 | 05:12 target changed to dummy, |
||
| 311 | 06:31 target changed back to Loop 2 10cm LH2, |
||
| 312 | 13:12 SHMS rotated to 6.76deg, shms_p=6.265,shms_theta=6.76,hms_p=-4.256,hms_theta=13.43, |
||
| 313 | |||
| 314 | 3116 production, 14:10-23:18 July 31st, target_type=Loop 2 10cm LH2,shms_p=6.265,shms_theta=6.76,hms_p=-4.256,hms_theta=13.43,I=40uA |
||
| 315 | Target changed to dummp at 14:35, then changed back to Loop 2 at 15:34, |
||
| 316 | 17:58 SHMS angle changed to 7.91 deg, shms_p=6.538,shms_theta=7.91,hms_p=-3.944,hms_theta=17.5,I=40-65uA |
||
| 317 | |||
| 318 | 3117 production, 23:18-09:24, target_type=Loop 2 10cm LH2,shms_p=6.538,shms_theta=7.91,hms_p=-3.944,hms_theta=17.5, I=40uA |
||
| 319 | Target changed to dummy at 01:40, then to Loop 1 10 cm LD2 at 02:40, beam stopped at 05:00 |
||
| 320 | |||
| 321 | 3118 production, 09:24-21:59, target_type=Loop 1 10cm LD2,shms_p=6.538,shms_theta=7.91,hms_p=-3.944,hms_theta=17.5, I=40uA |
||
| 322 | 3119 production, 21:59-11:40, target_type=Loop 1 10cm LD2,shms_p=6.538,shms_theta=7.91,hms_p=-3.944,hms_theta=17.5, I=40uA, |
||
| 323 | SHMS rotated to 9.915deg at 5:25 20220802, after that, shms_p=6.538,shms_theta=9.91,hms_p=-3.944,hms_theta=17.5, I=50-65uA |
||
| 324 | |||
| 325 | 3120 production,11:42-23:49, target_type=Loop 1 10cm LD2,shms_p=6.538,shms_theta=9.91,hms_p=-3.944,hms_theta=17.5 |
||
| 326 | Target changed to 10cm dummy at 17:39, then changed to Loop 2 10 cm LH2 at 18:47, this run finished at 23:49 |
||
| 327 | |||
| 328 | 3121 production, 23:51-08:56, target_type=Loop 2 10cm LH2,shms_p=6.538,shms_theta=9.91,hms_p=-3.944,hms_theta=17.5, I=70uA |
||
| 329 | no beam for the fisrt 8 hours |
||
| 330 | |||
| 331 | GEM gas bottle might be emtpy ... need to verify from data |
||
| 332 | |||
| 333 | 3122 production, 08:56-11:14, target_type=Loop 2 10cm LH2,shms_p=6.538,shms_theta=9.91,hms_p=-3.944,hms_theta=17.5, I=70uA |
||
| 334 | SHMS rotated to 11.91 deg at 10:25, shms_p=6.538,shms_theta=11.91,hms_p=-3.944,hms_theta=17.5, |
||
| 335 | |||
| 336 | 3123 production, 11:14-14:47, target_type=Loop 2 10cm LH2,shms_p=6.538,shms_theta=11.91,hms_p=-3.944,hms_theta=17.5, I=70uA |
||
| 337 | 3124 production, 14:48-17:39, target_type=Loop 2 10cm LH2,shms_p=6.538,shms_theta=11.91,hms_p=-3.944,hms_theta=17.5, I=70uA |
||
| 338 | |||
| 339 | GEM HV is off |
||
| 340 | |||
| 341 | 3125 production, 17:45-20:07, target_type=Loop 2 10cm LH2,shms_p=6.538,shms_theta=11.91,hms_p=-3.944,hms_theta=17.5, I=70uA |
||
| 342 | |||
| 343 | change target to dummy |
||
| 344 | 3126 production, 20:22-21:07, target_type=dummy,shms_p=6.538,shms_theta=11.91,hms_p=-3.944,hms_theta=17.5, I=70uA |
||
| 345 | |||
| 346 | ==================================================================== |
||
| 347 | FADC slot 7.11 DAC value changed from 3270 to 3170 |
||
| 348 | slot 7: |
||
| 349 | #FADC250_ALLCH_DAC 3205 3170 3170 3195 3195 3190 3200 3195 3180 3205 3190 3215 3170 3215 3175 3205 |
||
| 350 | #FADC250_ALLCH_DAC 3150 3120 3120 3150 3150 3140 3150 3150 3130 3160 3140 3270 3120 3170 3130 3160 |
||
| 351 | FADC250_ALLCH_DAC 3150 3120 3120 3150 3150 3140 3150 3150 3130 3160 3140 3170 3120 3170 3130 3160 |
||
| 352 | |||
| 353 | 3127 production, 20:22-21:07, target_type=dummy,shms_p=6.538,shms_theta=11.91,hms_p=-3.944,hms_theta=17.5, I=70uA |
||
| 354 | about 85k events only, but they are good events |
||
| 355 | |||
| 356 | changed target to Loop 1 |
||
| 357 | |||
| 358 | 3128 production, 21:11-22:41, target_type=Loop 1 10cm LD2,shms_p=6.538,shms_theta=11.91,hms_p=-3.944,hms_theta=17.5, I=70uA |
||
| 359 | 3129 production, 22:42-02:35, target_type=Loop 1 10cm LD2,shms_p=6.538,shms_theta=11.91,hms_p=-3.944,hms_theta=17.5, I=70uA |
||
| 360 | |||
| 361 | changed target to Loop 2 LH2, then changed target to Loop 2 LH2. |
||
| 362 | HMS was rotated 20.57deg, SHMS was rotated to 8.46deg. |
||
| 363 | |||
| 364 | |||
| 365 | 3130 production, 02:41-08:15, target_type=Loop 2 10cm LH2,shms_p=6.719,shms_theta=8.46,hms_p=-3.716,hms_theta=20.58, |
||
| 366 | 3131 production, 08:16-12:41, target_type=Loop 2 10cm LH2,shms_p=6.719,shms_theta=8.46,hms_p=-3.716,hms_theta=20.58, |
||
| 367 | 3132 production, 12:41-17:03, target_type=Loop 2 10cm LH2,shms_p=6.719,shms_theta=8.46,hms_p=-3.716,hms_theta=20.58, |
||
| 368 | no beam since 15:30, thunderstorems affect accelerator |
||
| 369 | |||
| 370 | changed target to dummy |
||
| 371 | |||
| 372 | 3133 no beam, pedestal run, 17:05-17:05 |
||
| 373 | 3134 no beam, pedestal run, 17:07-17:07 |
||
| 374 | 3135 production, 17:11-21:06, target_type=dummy,shms_p=6.719,shms_theta=8.46,hms_p=-3.716,hms_theta=20.58, |
||
| 375 | no beam for the first 100k events |
||
| 376 | 3136 production, 17:11-23:58, target_type=dummy,shms_p=6.719,shms_theta=8.46,hms_p=-3.716,hms_theta=20.58, I=40uA |
||
| 377 | no beam for the first 100k events |
||
| 378 | 3137 production, 23:58-0:16, target_type=dummy,shms_p=6.719,shms_theta=8.46,hms_p=-3.716,hms_theta=20.58, I=40uA |
||
| 379 | SHMS rotated to 10.475deg at 0:08 |
||
| 380 | 3138 production, 0:16-1:09, target_type=dummy,shms_p=6.719,shms_theta=10.475,hms_p=-3.716,hms_theta=20.58, I=40uA |
||
| 381 | 3139 production, 1:10-02:11, target_type=dummy,shms_p=6.719,shms_theta=10.475,hms_p=-3.716,hms_theta=20.58, I=40uA |
||
| 382 | |||
| 383 | changed target to Loop 2 10cm LH2 |
||
| 384 | |||
| 385 | 3140 production, 2:13-8:25, target_type=Loop 2 10cm LH2,shms_p=6.719,shms_theta=10.475,hms_p=-3.716,hms_theta=20.58, I=40uA |
||
| 386 | 3141 no beam, pedestal run |
||
| 387 | |||
| 388 | 08:43 HV is off |
||
| 389 | 10:23 HV is ON |
||
| 390 | |||
| 391 | 3142 no beam, pedestal run |
||
| 392 | 3143 no beam, pedestal run |
||
| 393 | |||
| 394 | 3144 production, 14:08-21:39, target_type=Loop 2 10cm LH2,shms_p=6.719,shms_theta=10.475,hms_p=-3.716,hms_theta=20.58, I=40uA |
||
| 395 | no beam during the first 100k events. Beam was on for just several minutes. |
||
| 396 | 3145 production, 21:40-11:20, target_type=Loop 2 10cm LH2,shms_p=6.719,shms_theta=10.475,hms_p=-3.716,hms_theta=20.58, I=70uA |
||
| 397 | no beam during the first 100k events. beam started at 4:05 |
||
| 398 | 3146 production, 11:20-15:44, target_type=Loop 2 10cm LH2,shms_p=6.719,shms_theta=10.475,hms_p=-3.716,hms_theta=20.58, I=70uA |
||
| 399 | SHMS angle changed to 12.49 deg at 12:16 |
||
| 400 | |||
| 401 | 3147 production, 15:44-01:39 August 7th, target_type=Loop 2 10cm LH2,shms_p=6.719,shms_theta=12.49,hms_p=-3.716,hms_theta=20.58 |
||
| 402 | daq rate 1.5kHz @ I=70uA |
||
| 403 | 18:20 target changed to dummy |
||
| 404 | 22:00 target changed to Loop 2 LH2, I=70uA |
||
| 405 | |||
| 406 | |||
| 407 | 3148 production, 01:39-07:24 August 7th, target_type=Loop 2 10cm LH2,shms_p=6.719,shms_theta=12.49,hms_p=-3.716,hms_theta=20.58 |
||
| 408 | beam stopped at 7:00 |
||
| 409 | |||
| 410 | 7:30 SHMS rotated 12.28 deg, HMS rotated to 15.06 deg, target was changed to Loop 1 10cm LD2 |
||
| 411 | |||
| 412 | =================================================================== |
||
| 413 | FADC250_W_WIDTH = 240 for both slot7 and slot8 |
||
| 414 | |||
| 415 | 3149 pedestal run, no beam |
||
| 416 | 3150 pedestal run, no beam |
||
| 417 | 3151 production, 07:34-13:32, target_type=Loop 1 10cm LD2, |
||
| 418 | shms_p=-5.127,shms_theta=12.29,hms_p=-5.309,hms_theta=15.04, I=75uA |
||
| 419 | From 10:47 to 11:11, target was changed to dummy (I=40uA), at |
||
| 420 | 11:42 target was changed back to Loop 1 10cm LD2, I=75uA |
||
| 421 | At 12:42, SHMS was ratated to 14.29deg, at this angle, daq rate 5.5kHz |
||
| 422 | subrun 0-50: SHMS=12.29deg |
||
| 423 | subrun 51: SHMS=12.29deg and 14.29deg |
||
| 424 | subrun 52-60: SHMS=14.29deg |
||
| 425 | |||
| 426 | 3152 production, 13:32-17:50 August 7th, target_type=Loop 1 10cm LD2, |
||
| 427 | shms_p=-5.127,shms_theta=14.29,hms_p=-5.309,hms_theta=15.04, |
||
| 428 | I=75uA, daq rate 5.5kHz |
||
| 429 | 16:37 target changed to dummy, I=40uA |
||
| 430 | 17:25 target changed to Loop 1 10cm LD2, I=75uA |
||
| 431 | 17:42 SHMS rotated to 16.28deg, I=75uA, daq rate 5.5kHz |
||
| 432 | |||
| 433 | subrun 0-42: SHMS=14.29deg |
||
| 434 | subrun 43: SHMS=14.29deg and 16.28deg |
||
| 435 | subrun 44-45: SHMS=16.28deg |
||
| 436 | |||
| 437 | |||
| 438 | 3153 production, 17:50-19:26, target_type=Loop 1 10cm LD2, shms_p=-5.127,shms_theta=16.28,hms_p=-5.309,hms_theta=15.04,I=75uA |
||
| 439 | |||
| 440 | 3154 production, 19:26-00:01, target_type=Loop 1 10cm LD2, shms_p=-5.127,shms_theta=16.28,hms_p=-5.309,hms_theta=15.04,I=75uA |
||
| 441 | 23:35 changed to dummy target,I=40uA, daq rate 1.6kHz |
||
| 442 | |||
| 443 | 3155 production, 00:02-00:33, target_type=dummy, shms_p=-5.127,shms_theta=16.28,hms_p=-5.309,hms_theta=15.04,I=40uA, |
||
| 444 | daq rate 1.6kHz, beam stopped at 00:14 |
||
| 445 | |||
| 446 | 20220808 00:14, rotated SHMS to 16.91 deg, HMS to 19.92 deg |
||
| 447 | |||
| 448 | 3156 production, 00:37-01:48, target_type=dummy,shms_p=-5.512,shms_theta=16.94,hms_p=-4.752,hms_theta=19.91,I=40uA, |
||
| 449 | daq rate 1.6kHz, |
||
| 450 | target changed to Loop 1 10cm LD2 at 01:16, |
||
| 451 | target_type=Loop 1 10cm LD2,shms_p=-5.512,shms_theta=16.94,hms_p=-4.752,hms_theta=19.91, I=75uA |
||
| 452 | |||
| 453 | 3157 production, 00:37-03:11, target_type=Loop 1 10cm LD2,shms_p=-5.512,shms_theta=16.94,hms_p=-4.752,hms_theta=19.91, I=75uA |
||
| 454 | daq rate 5.6kHz |
||
| 455 | |||
| 456 | 3158 production, 03:11-09:11, target_type=Loop 1 10cm LD2,shms_p=-5.512,shms_theta=16.94,hms_p=-4.752,hms_theta=19.91, I=75uA |
||
| 457 | no beam for the fisrt 100k events, daq rate 5.6kHz |
||
| 458 | |||
| 459 | 3159 production, 09:11-10:09, target_type=Loop 1 10cm LD2,shms_p=-5.512,shms_theta=16.94,hms_p=-4.752,hms_theta=19.91, I=75uA |
||
| 460 | daq rate 5.6kHz |
||
| 461 | |||
| 462 | 3160 production, 10:10-13:24, target_type=Loop 1 10cm LD2,shms_p=-5.512,shms_theta=16.94,hms_p=-4.752,hms_theta=19.91, I=75uA |
||
| 463 | daq rate 5.6kHz |
||
| 464 | |||
| 465 | 3161 production, 13:24-15:44, target_type=Loop 1 10cm LD2,shms_p=-5.512,shms_theta=16.94,hms_p=-4.752,hms_theta=19.91, I=75uA |
||
| 466 | daq rate 5.6kHz |
||
| 467 | target changed to dummy at 15:18, I=40uA |
||
| 468 | |||
| 469 | 3162 production, 15:44-16:32, target_type=dummy,shms_p=-5.512,shms_theta=16.94,hms_p=-4.752,hms_theta=19.91, I=40uA |
||
| 470 | beam stop at 16:03 to prepare for optic |
||
| 471 | |||
| 472 | 16:09 target changed to single foil Carbon 0.5%, SHMS rotated to 6.8deg and HMS ratated to 12.5deg |
||
| 473 | |||
| 474 | 3163 production, 16:32-17:22, target_type=Carbon 0.5%, shms_p=-8.55,shms_theta=6.8,hms_p=-5.587,hms_theta=12.48,I=75uA |
||
| 475 | no beam for the first 100k events, daq rate 160Hz |
||
| 476 | |||
| 477 | Target move to Optics 1 +/- 8cm, 2 Foils Carbon |
||
| 478 | |||
| 479 | 3164 production, 17:22-17:47, target_type=Optics 1 z=+/-8cm, 2 Foil, shms_p=-8.55,shms_theta=6.8,hms_p=-5.587,hms_theta=12.48,I=50uA |
||
| 480 | no beam for the first 100k events, daq rate 80Hz |
||
| 481 | |||
| 482 | 3165 production, 17:47-18:02, target_type=Optics 1 z=+/-8cm, 2 Foil, shms_p=-8.55,shms_theta=6.8,hms_p=-5.587,hms_theta=12.48,I=50uA |
||
| 483 | only 30k triggers |
||
| 484 | |||
| 485 | 18:00 target changed to Loop 2 10cm LH2 |
||
| 486 | |||
| 487 | 3166 production, 18:05-18:57, target_type=Loop2 10cm LH2, shms_p=-8.55,shms_theta=6.8,hms_p=-5.587,hms_theta=12.48, |
||
| 488 | I=60uA (18:06), then 20uA(18:08) then 10uA(18:10) |
||
| 489 | daq rate 130Hz @20uA, 65Hz @10uA |
||
| 490 | SHMS rotated to 7.5deg at 18:28, shms_p=-8.55,shms_theta=7.5,hms_p=-5.587,hms_theta=12.48,I=15uA |
||
| 491 | |||
| 492 | 18:57 Target changet to Optics 1 +/- 8 cm |
||
| 493 | 3167 production, 18:59-19:37, target_type=Optics 1 z=+/-8cm, shms_p=-8.55,shms_theta=7.5,hms_p=-5.587,hms_theta=12.48,I=50uA |
||
| 494 | daq rate 90Hz @50uA |
||
| 495 | target changed to Carbon 0.5% at 19:20, I=75uA |
||
| 496 | |||
| 497 | 19:38 SHMS rotated to 8.3 deg |
||
| 498 | |||
| 499 | 3168 production, 19:39-19:54, target_type=Carbon 0.5%, shms_p=-8.55,shms_theta=8.3,hms_p=-5.587,hms_theta=12.48,I=75uA |
||
| 500 | |||
| 501 | 19:54 Target changet to Optics 1 +/- 8 cm |
||
| 502 | |||
| 503 | 3169 production, 19:54-20:34, target_type=Optics 1 z=+/-8cm, shms_p=-8.55,shms_theta=8.3,hms_p=-5.587,hms_theta=12.48,I=50uA |
||
| 504 | no beam for the first 100k events, |
||
| 505 | daq rate 90Hz @50uA |
||
| 506 | 20:24 target changed to Loop 2 10cm LH2, beam came back at 20:29 |
||
| 507 | |||
| 508 | 3170 production, 20:34-21:23, target_type=Loop 2 10cm LH2, shms_p=-8.55,shms_theta=8.3,hms_p=-5.587,hms_theta=12.48,I=20uA |
||
| 509 | daq rate 160Hz @20uA, beam stopped at 21:06 |
||
| 510 | |||
| 511 | =================================================== |
||
| 512 | HV map for FTBF beam test |
||
| 513 | |||
| 514 | HV for FADC channel 0 to 7: |
||
| 515 | # # SDU#4 SDU#5 THU#2 SDU#2 SDU#1 THU#1 |
||
| 516 | SC1 SC2 Ecal1 Ecal2 Ecal3 PrSh1 PrSh2 PrSh3 |
||
| 517 | -1260 -1280 -760 -860 -880 -1640 -1480 -1800 <<== 1~8GeV |
||
| 518 | |||
| 519 | This setting was also used for some 1GeV calibrations |
||
| 520 | -1260 -1300 -880 -920 -940 -1640 -1440 -1800 <<== <1GeV |
||
| 521 | |||
| 522 | --------------------------------------- |
||
| 523 | 4GeV calibration start |
||
| 524 | 1009 Moved table to (1505, 337), center on the left block 51,563 events |
||
| 525 | 1010 Moved table to (1389, 337), center on the right block 77621 events |
||
| 526 | 1011 Moved table to (1447, 243), center on the top block 52034 events |
||
| 527 | 4GeV calibration end |
||
| 528 | --------------------------------------- |
||
| 529 | |||
| 530 | |||
| 531 | --------------------------------------- |
||
| 532 | 1GeV calibration start, MWPC 1 is not working |
||
| 533 | 796 move to Top block, use this New HV: 1260, 1300, 880, 920, 940, 1640, 1440, 1800, FADC ch11 and ch12 swapped. |
||
| 534 | These two channels are down stream Cherenkov. After swapped, Down cherenkov inner PMT is in ch12. |
||
| 535 | 797 |
||
| 536 | |||
| 537 | 798 move to left block, point B (1498,347) |
||
| 538 | 799 |
||
| 539 | 800 beam on the center of the left block, point B |
||
| 540 | 801 |
||
| 541 | 802 beam on the center of the right block, point C (1382,347) |
||
| 542 | |||
| 543 | 803 beam on the center of the top block, point A (1440, 252) |
||
| 544 | 1GeV calibration end |
||
| 545 | --------------------------------------- |
||
| 546 | ===================================================== |
||
| 547 | |||
| 548 | ----------------------------------------------- |
||
| 549 | change HV to the same as FermiLab |
||
| 550 | # # SDU#4 SDU#5 THU#2 SDU#2 SDU#1 THU#1 |
||
| 551 | SC1 SC2 Ecal1 Ecal2 Ecal3 PrSh1 PrSh2 PrSh3 |
||
| 552 | -1260 -1280 -760 -860 -880 -1640 -1480 -1800 <<== 1~8GeV |
||
| 553 | ===> |
||
| 554 | |||
| 555 | slot10.7(SDU#1) slot10.8(THU#1) slot10.9(SDU#2) |
||
| 556 | -1480 -1800 -1640 |
||
| 557 | |||
| 558 | slot2.0(SDU#4) slot2.1(THU#2) slot2.2(SDU#5) |
||
| 559 | -760 -880 -860 |
||
| 560 | ----------------------------------------------- |
||
| 561 | |||
| 562 | 21:40 SHMS rotated to 12.25deg, HMS rotated to 15.7deg |
||
| 563 | |||
| 564 | ----------------------------------------------- |
||
| 565 | |||
| 566 | 22:11 target changed to dummy |
||
| 567 | |||
| 568 | 3171 production, 22:10-22:21,target_type=dummy,shms_p=-8.035,shms_theta=12.25,hms_p=-6.792,hms_theta=15.69,I=40uA |
||
| 569 | no beam at the first 100k triggers |
||
| 570 | |||
| 571 | 22:34 SHMS rotated to 13.15deg, HMS rotated to 17.72deg |
||
| 572 | |||
| 573 | 3172 production, 22:22-23:22,target_type=dummy,shms_p=-8.035,shms_theta=13.15,hms_p=-6.792,hms_theta=17.72, |
||
| 574 | no beam at the first 100k triggers, beam came at 22:32. only 9 minutes of dummy data |
||
| 575 | 22:41 stopped beam to changed target to Loop 2 10cm LH2 |
||
| 576 | 22:46 beam came back and then went away at 22:53 due to HMS Q1 interlock, |
||
| 577 | 23:12 beam came back again, shms_p=-8.035,shms_theta=13.15,hms_p=-6.792,hms_theta=17.72, I=60uA, |
||
| 578 | daq rate 1.1kHz, 10 minutes of LH2 data |
||
| 579 | |||
| 580 | 3173 production, 23:22-23:52 August 8th, |
||
| 581 | target=Loop 2 10cm LH2, shms_p=-8.035,shms_theta=13.15,hms_p=-6.792,hms_theta=17.72, I=60uA |
||
| 582 | daq rate 1.1kHz |
||
| 583 | beam stopped at 23:38 then SHMS rotated to 14.02 deg and HMS rotated to 19.86deg |
||
| 584 | beam came back at 23:49, I=75uA, shms_p=-8.035,shms_theta=14.02,hms_p=-6.792,hms_theta=19.86, |
||
| 585 | |||
| 586 | subrun 0: SHMS=13.15deg |
||
| 587 | subrun 1: SHMS=13.15deg and 14.02deg |
||
| 588 | |||
| 589 | |||
| 590 | 3174 production, 23:52-02:03 August 9th, |
||
| 591 | target=Loop 2 10cm LH2, shms_p=-8.035,shms_theta=14.02,hms_p=-6.792,hms_theta=19.86, I=75uA |
||
| 592 | daq rate 1.9kHz |
||
| 593 | 0:51 target changed to dummy, I=40uA, |
||
| 594 | 01:16 SHMS rotated to 14.895deg, shms_p=-8.035,shms_theta=14.9,hms_p=-6.792,hms_theta=19.86, I=75uA |
||
| 595 | 01:45 target changed to Loop 2 LH2, beam came back at 01:49 but stopped at 01:58 |
||
| 596 | subrun 0-3 are for 14.02 deg SHMS |
||
| 597 | subrun 4 are for both 14.02 and 14.9 deg SHMS |
||
| 598 | subrun 5-6 are for 14.9 deg SHMS |
||
| 599 | |||
| 600 | |||
| 601 | 3175 production, 02:03-08:35,target=Loop 2 10cm LH2, shms_p=-8.035,shms_theta=14.9,hms_p=-6.792,hms_theta=19.86,I=75uA |
||
| 602 | no beam at the first 100k triggers,beam came back at 03:18, |
||
| 603 | 04:45, SHMS rotated to 15.785deg, shms_p=-8.035,shms_theta=15.79,hms_p=-6.792,hms_theta=19.86,I=70uA |
||
| 604 | 06:05, target changed to dummy, I=40uA, took 8 minutes of data |
||
| 605 | 06:25, SHMS rotated to 13.58deg, took 8 minutes of data for dummy, I=40uA, |
||
| 606 | 06:47, changed target to Loop2 LH2, shms_p=-7.913,shms_theta=13.58,hms_p=-6.792,hms_theta=19.86, |
||
| 607 | 08:00, SHMS rotated to 12.69deg, took 4 minutes of data till beam stopped at 08:04 to change pass, |
||
| 608 | shms_p=-8.035,shms_theta=12.69,hms_p=-6.792,hms_theta=16.03,I=70uA |
||
| 609 | |||
| 610 | ============================================================================================================== |
||
| 611 | Aug 9th 08:05 Accelerator changed beam from 5-pass to 4-pass for Hall C. Beam energy is 8.494GeV |
||
| 612 | |||
| 613 | 3176 pedestal, 100k pulse triggers, no beam |
||
| 614 | 3177 pedestal, 100k pulse triggers, no beam |
||
| 615 | |||
| 616 | 08:48 HMS rotated to 20.905deg, SHMS rotated to 6.93deg |
||
| 617 | |||
| 618 | |||
| 619 | 3178 production, 09:02-15:08, target=Loop 2 LH2, shms_p=6.265,shms_theta=6.93,hms_p=-2.193,hms_theta=20.89, |
||
| 620 | no beam at the first 100k triggers,beam came back at 14:02, |
||
| 621 | 14:00 target changed to Carbon Hole, I=5uA for 6 minutes, |
||
| 622 | 14:18 target changed to Loop 2 LH2, I=35uA, beam came back at 14:20, daq rate 220Hz, beam stopped at 14:46 |
||
| 623 | |||
| 624 | 3179 production, 15:09-21:31, target=Loop 2 10cm LH2, shms_p=6.265,shms_theta=6.93,hms_p=-2.193,hms_theta=20.89, |
||
| 625 | no beam at the first 100k triggers,beam came back at 17:31, I=60uA |
||
| 626 | |||
| 627 | 22:37 target changed to dummy |
||
| 628 | |||
| 629 | 3180 production, 21:31-22:28, target=dummy, shms_p=6.265,shms_theta=6.93,hms_p=-2.193,hms_theta=20.89 |
||
| 630 | no beam for the first 100k triggers, |
||
| 631 | 21:35 beam arrived, beam current linear ran up from 3uA to 23uA at 21:55 to do Ion Chamber calibration |
||
| 632 | From 21:56, I=40UA till 22:15 |
||
| 633 | 22:15 SHMS rotated to 8.93 deg, beam came back at 22:20 |
||
| 634 | target=dummy, shms_p=6.265,shms_theta=8.93,hms_p=-2.193,hms_theta=20.89 |
||
| 635 | |||
| 636 | 3181 production, 22:28-22:44, target=dummy, shms_p=6.265,shms_theta=8.93,hms_p=-2.193,hms_theta=20.89 |
||
| 637 | |||
| 638 | 22:46 target changed to Loop 2 LH2 |
||
| 639 | |||
| 640 | 3182 production, 22:48-00:05 August 10th, |
||
| 641 | target=Loop 2 10cm LH2, shms_p=6.265,shms_theta=8.93,hms_p=-2.193,hms_theta=20.89, I=80uA |
||
| 642 | no beam for the first 100k triggers,daq rate 770Hz |
||
| 643 | |||
| 644 | 3183 production, 00:05-01:22, |
||
| 645 | target=Loop 2 10cm LH2, shms_p=6.265,shms_theta=8.93,hms_p=-2.193,hms_theta=20.89, I=80uA |
||
| 646 | daq rate 770Hz, beam stopped at 00:49 due to Vacuum issue at VBV3H08A |
||
| 647 | |||
| 648 | 3184 pedestal, pulse run, 01:23-10:26, no beam for the whole run,1000006 events |
||
| 649 | 3185 pedestal, pulse run, 11:29-11:29, no beam |
||
| 650 | ========================================================================================= |
||
| 651 | Add "PreSh_t && Shower_t" into TS#2, 20220810 |
||
| 652 | |||
| 653 | 3186 pedestal, pulse run, 14:54-15:19, no beam |
||
| 654 | 3187 pedestal, pulse run, 15:47-15:47, no beam |
||
| 655 | |||
| 656 | 16:00 SHMS rotated to 5.5 deg. |
||
| 657 | |||
| 658 | 3188 production, 17:21-23:07, target=Loop 2 10cm LH2, shms_p=6.265,shms_theta=5.53,hms_p=-2.193,hms_theta=20.89, |
||
| 659 | no beam for the sirst 100k triggers, tune beam arrived at 17:52, real beam arrived at 21:34 |
||
| 660 | |||
| 661 | 23:09 Changed FADC slot8.9 DAC value from 3130 to 3440. After feed it into FIFO, its pedestal changed |
||
| 662 | |||
| 663 | 3189 production,23:10-02:38 August 11th, |
||
| 664 | target=Loop 2 LH2, shms_p=6.265,shms_theta=5.53,hms_p=-2.193,hms_theta=20.89,I=30uA |
||
| 665 | daq rate 320Hz |
||
| 666 | |||
| 667 | 3190 production,02:38-04:38, target=Loop 2 LH2, shms_p=6.265,shms_theta=5.53,hms_p=-2.193,hms_theta=20.89,I=30uA |
||
| 668 | daq rate 320Hz |
||
| 669 | 03:05 target changed to dummy, I=40uA |
||
| 670 | 03:45 HMS rotated to 28.41deg, |
||
| 671 | target=dummy,shms_p=6.538,shms_theta=5.53,hms_p=-1.882,hms_theta=28.41, I=40uA, daq rate 320Hz |
||
| 672 | |||
| 673 | 3191 production, 04:39-09:03, target=dummy,shms_p=6.538,shms_theta=5.53,hms_p=-1.882,hms_theta=28.41, |
||
| 674 | I=40uA, daq rate 320Hz |
||
| 675 | 4:50 target chagned to Loop 2 LH2, I=35uA |
||
| 676 | target=Loop 2 LH2, shms_p=6.265,shms_theta=5.53,hms_p=-2.193,hms_theta=28.41,I=35uA |
||
| 677 | Acceleration maintenence started from 6:30 |
||
| 678 | |||
| 679 | 3192 production, 19:55-01:52 August 12th, target=Loop 2 LH2,shms_p=6.538,shms_theta=5.53,hms_p=-1.882,hms_theta=28.41, |
||
| 680 | no beam for the first 100k triggers, tune beam came back at 23:32 |
||
| 681 | |||
| 682 | 3193 production, 01:52-09:28, target=Loop 2 LH2,shms_p=6.538,shms_theta=5.53,hms_p=-1.882,hms_theta=28.41, |
||
| 683 | daq rate = 400Hz @ I=38uA. Beam stopped at 09:10 |
||
| 684 | |||
| 685 | 09:21 A beam energy measurement is scheduled at 9:30-ish AM. Target parked at 'home'. |
||
| 686 | 11:29 Target changed to Loop 2 LH2 |
||
| 687 | |||
| 688 | 3194 production, 11:35-12:26, target=Loop 2 LH2,shms_p=6.538,shms_theta=5.53,hms_p=-1.882,hms_theta=28.41, |
||
| 689 | no beam for the first 100k triggers, beam came back at 11:50 and went away at 12:01 |
||
| 690 | |||
| 691 | 3195 production, 14:05-14:41, target=Loop 2 LH2,shms_p=6.538,shms_theta=5.53,hms_p=-1.882,hms_theta=28.41, |
||
| 692 | daq rate = 400Hz @ I=38uA. |
||
| 693 | |||
| 694 | 3196 production, 14:42-16:46, target=Loop 2 LH2,shms_p=6.538,shms_theta=5.53,hms_p=-1.882,hms_theta=28.41, |
||
| 695 | no beam for the first 100k triggers, beam came back at 14:49 |
||
| 696 | beam stopped at 16:12 |
||
| 697 | |||
| 698 | 3197 production, 16:49-00:54 August 13th, target=Loop 2 LH2, |
||
| 699 | shms_p=6.538,shms_theta=5.53,hms_p=-1.882,hms_theta=28.41, |
||
| 700 | no beam for the first 100k triggers, beam came back at 17:48 |
||
| 701 | |||
| 702 | |||
| 703 | 3198 production, 00:54-02:28, target=Loop 2 LH2,shms_p=6.538,shms_theta=5.53,hms_p=-1.882,hms_theta=28.41, |
||
| 704 | daq rate = 450Hz @ I=40uA. |
||
| 705 | 01:09 SHMS rotated to 7.47 deg, trying to ramp beam current up to 80uA |
||
| 706 | subrun 0: 5.53deg and 7.47deg |
||
| 707 | subrun 1-3: 7.47deg |
||
| 708 | |||
| 709 | |||
| 710 | 3199 production, 02:28-12:05, target=Loop 2 LH2,shms_p=6.538,shms_theta=7.47,hms_p=-1.882,hms_theta=28.41, |
||
| 711 | daq rate = 1024Hz @ I=79uA. |
||
| 712 | |||
| 713 | 3200 production, 12:05-14:42, target=Loop 2 LH2,shms_p=6.538,shms_theta=7.47,hms_p=-1.882,hms_theta=28.41, |
||
| 714 | daq rate = 1024Hz @ I=79uA. |
||
| 715 | 14:28 target changed to dummy, I=40uA, daq rate 320Hz |
||
| 716 | |||
| 717 | 3201 production, 14:42-16:01, target=dummy,shms_p=6.538,shms_theta=7.47,hms_p=-1.882,hms_theta=28.41, |
||
| 718 | I=40uA, daq rate 320Hz |
||
| 719 | |||
| 720 | 16:01 SHMS rotated to 9.5deg, target changed to Loop 2 LH2 |
||
| 721 | |||
| 722 | 3202 production, 14:01-17:29, target=Loop 2 LH2,shms_p=6.538,shms_theta=9.5,hms_p=-1.882,hms_theta=28.41, |
||
| 723 | no beam for the first 100k triggers, beam came back 16:04 |
||
| 724 | daq rate = 1024Hz @ I=79uA. |
||
| 725 | beam stopped at 17:13, 4th pass separation issue |
||
| 726 | |||
| 727 | 3203 production, 17:37-00:35 August 14th, target=Loop 2 LH2,shms_p=6.538,shms_theta=9.5,hms_p=-1.882,hms_theta=28.41, |
||
| 728 | no beam for the first 100k triggers, beam came back 17:41 |
||
| 729 | daq rate 1.5kHz @ I=80uA |
||
| 730 | |||
| 731 | 3204 production, 00:35-03:53 August 14th, target=Loop 2 LH2,shms_p=6.538,shms_theta=9.5,hms_p=-1.882,hms_theta=28.41, |
||
| 732 | daq rate 1.5kHz @ I=80uA |
||
| 733 | 03:31 target changed to dummy, I=40uA |
||
| 734 | |||
| 735 | 3205 production, 03:53-04:13, target=dummy,shms_p=6.538,shms_theta=9.5,hms_p=-1.882,hms_theta=28.41, |
||
| 736 | I=40uA, daq rate 512Hz |
||
| 737 | |||
| 738 | 3206 production, 04:13-04:31, target=dummy,shms_p=6.538,shms_theta=9.5,hms_p=-1.882,hms_theta=28.41, |
||
| 739 | I=40uA, daq rate 512Hz |
||
| 740 | |||
| 741 | 04:38 target changed to Loop 2 LH2 |
||
| 742 | 04:50 SHMS rotated to 8.93 deg, HMS rotated to 20.89deg |
||
| 743 | |||
| 744 | 3207 production, 05:01-13:02, target=Loop 2 LH2,shms_p=6.265,shms_theta=8.93,hms_p=-2.193,hms_theta=20.89, |
||
| 745 | no beam for the first 100k triggers, beam came back 7:35, I=80uA |
||
| 746 | beam stopped at 10:32, then came back at 12:15 |
||
| 747 | ~11:00 SHMS rotated to 30.02deg, HMS rotated to 19.56deg |
||
| 748 | target=Loop 2 LH2,shms_p=3.731,shms_theta=30.02,hms_p=-5.587,hms_theta=19.56 |
||
| 749 | ---> SHMS=8.93deg for subrun 0-3, SHMS=30.02deg for subrun 5 to 10, subrun 4 contains both angles <--- |
||
| 750 | |||
| 751 | 3208 production, 13:02-13:14, target=Loop 2 LH2,shms_p=3.731,shms_theta=30.02,hms_p=-5.587,hms_theta=19.56,I=80uA |
||
| 752 | At 13:07, target changed to dummy, I=40uA. Only 2 minutes of LH2 target then 2 minutes of dummy target |
||
| 753 | |||
| 754 | 3209 production, 13:14-13:25, target=dummy,shms_p=3.731,shms_theta=30.02,hms_p=-5.587,hms_theta=19.56, |
||
| 755 | daq rate 1.5kHz @ I=40uA |
||
| 756 | |||
| 757 | 3210 production, 13:25-13:30, target=dummy,shms_p=3.731,shms_theta=30.02,hms_p=-5.587,hms_theta=19.56, |
||
| 758 | daq rate 1.5kHz @ I=40uA, only 2 minutes of data then beam stopped for target change |
||
| 759 | |||
| 760 | 13:32 target changed to Loop 2 LH2, |
||
| 761 | 13:32 SHMS rotated to 28.355deg. HMS rotated to 20.905deg |
||
| 762 | |||
| 763 | 3211 production, 13:33-13:44, target=Loop 2 LH2, shms_p=3.996,shms_theta=28.40,hms_p=-5.321,hms_theta=20.91, |
||
| 764 | daq rate 1.8kHz @ I=80uA, beam stopped at 13:41 |
||
| 765 | |||
| 766 | 3212 production, 13:45-14:39, target=Loop 2 LH2, shms_p=3.996,shms_theta=28.40,hms_p=-5.321,hms_theta=20.91, |
||
| 767 | daq rate 1.8kHz @ I=80uA, |
||
| 768 | I stopped coda at 13:44 for previous run because HallC coda was stopped at 13:41 in order to adjust |
||
| 769 | SHMS angle from 28.44 to 28.355. |
||
| 770 | I was taking data during this adjustment. This change should not affect the ecal setup. |
||
| 771 | |||
| 772 | 14:40 target changed to dummy |
||
| 773 | |||
| 774 | 3213 production, 14:41-14:58. target=dummy, shms_p=3.996,shms_theta=28.355,hms_p=-5.321,hms_theta=20.91, |
||
| 775 | I=40uA, daq rate 620Hz |
||
| 776 | |||
| 777 | |||
| 778 | 15:00 target changed to Loop 2 LH2, |
||
| 779 | 15:30 Set HMS angle to 15.065 and SHMS angle to 9.455 deg, but stayed at 15.05 deg and 9.53deg |
||
| 780 | |||
| 781 | 3214 production, 15:01-17:03, target=Loop 2 LH2, shms_p=-7.559,shms_theta=9.53,hms_p=-6.477,hms_theta=15.05, |
||
| 782 | no beam for the first 100k triggers, beam came back 15:37, I=20uA |
||
| 783 | 15:51 target changed to dummy,shms_p=-7.559,shms_theta=9.53,hms_p=-6.477,hms_theta=15.05, |
||
| 784 | 16:38 SHMS rotated to 11.97 deg,HMS rotated to 15.62 deg |
||
| 785 | target=dummy, shms_p=-7.106,shms_theta=11.97,hms_p=-5.878,hms_theta=15.62, |
||
| 786 | 17:02 target changed to Loop 2 LH2, shms_p=-7.106,shms_theta=11.97,hms_p=-5.878,hms_theta=15.62, |
||
| 787 | subrun 0: SHMS=9.455deg and 11.97deg |
||
| 788 | subrun 1: SHMS=11.97deg |
||
| 789 | |||
| 790 | 3215 production, 17:05-17:50, target=Loop 2 LH2, shms_p=-7.106,shms_theta=11.97,hms_p=-5.878,hms_theta=15.62, |
||
| 791 | no beam for the first 100k triggers, beam came back at 15:11, I=10uA, |
||
| 792 | beam went away at 17:24 due to injector problem, only 3~4 minutes of beam in this run, 175k events |
||
| 793 | |||
| 794 | 3216 production, 17:57-19:49, target=Loop 2 LH2, shms_p=-7.106,shms_theta=11.97,hms_p=-5.878,hms_theta=15.62, |
||
| 795 | no beam for the first 100k triggers, beam came back at 18:17, I=10uA till 18:45 |
||
| 796 | 18:47 SHMS rotated to 13.92, HMS rotated to 17.96deg, beam came at 19:13, I=55uA |
||
| 797 | target=Loop 2 LH2, shms_p=-6.72,shms_theta=13.92,hms_p=-5.878,hms_theta=17.96, |
||
| 798 | subrun 0: SHMS=11.97deg and 13.92deg |
||
| 799 | subrun 1-4: SHMS=13.92deg |
||
| 800 | |||
| 801 | 3217 production, 19:49-19:54, target=Loop 2 LH2, shms_p=-6.72,shms_theta=13.92,hms_p=-5.878,hms_theta=17.96, |
||
| 802 | beam stopped at 19:51, only 256k events |
||
| 803 | |||
| 804 | 19:56 target changed to dummy |
||
| 805 | |||
| 806 | 3218 production, 19:57-20:11, target=dummy, shms_p=-6.72,shms_theta=13.92,hms_p=-5.878,hms_theta=17.96, |
||
| 807 | |||
| 808 | 20:12 SHMS rotated to 14.82 deg, HMS rotated to 20.37deg |
||
| 809 | |||
| 810 | 3219 production, 20:13-20:36, target=dummy, shms_p=-6.538,shms_theta=14.82,hms_p=-5.878,hms_theta=20.37, |
||
| 811 | no beam for the first 100k triggers, beam came back at 20:20, I=40uA, daq rate 1920Hz |
||
| 812 | |||
| 813 | 20:35 target changed to Loop 2 LH2 |
||
| 814 | |||
| 815 | 3220 production, 20:40:21:16, target=Loop 2 LH2, shms_p=-6.538,shms_theta=14.82,hms_p=-5.878,hms_theta=20.37, |
||
| 816 | no beam for the first 100k triggers, beam came back at 20:42, I=80uA, |
||
| 817 | beam stopped at 21:12 to prepare optics runs |
||
| 818 | |||
| 819 | 21:18 HMS rotated to 13 deg, target changed to Carbon 0.5% |
||
| 820 | |||
| 821 | 3221 production, 21:16-22:05, target=Carbon 0.5%, |
||
| 822 | Sieve Optics,shms_p=-6.538,shms_theta=14.82,hms_p=-5.878,hms_theta=12.98, |
||
| 823 | no beam for the first 100k triggers, beam came back at 21:30, |
||
| 824 | I=4.5uA --> daq rate 64Hz, |
||
| 825 | I=13uA --> daq rate 192Hz, |
||
| 826 | I=80uA --> daq rate 1216Hz, |
||
| 827 | 22:02 target changed to optics 1, Carbon +/- 8cm, I=3uA |
||
| 828 | |||
| 829 | |||
| 830 | 3222 production, 22:05-23:21, target=Opitcs 1, Carbon 2-foils +/- 8cm, |
||
| 831 | Sieve Optics,shms_p=-6.538,shms_theta=14.82,hms_p=-5.878,hms_theta=12.98, |
||
| 832 | ramp currunet from 2uA till 40 uA |
||
| 833 | I= 10uA --> daq rate 128Hz |
||
| 834 | I= 15uA --> daq rate 192Hz |
||
| 835 | I= 20uA --> daq rate 256Hz |
||
| 836 | I= 49.5uA --> daq rate 576Hz |
||
| 837 | |||
| 838 | 23:24 target changed to Loop LH2 |
||
| 839 | 23:30 SHMS rotated to 16.100deg, HMS rotated to 17.235deg |
||
| 840 | |||
| 841 | 3223 production, 23:27-00:06, target=Loop 2 LH2, shms_p=-6,shms_theta=16.17,hms_p=-5.587,hms_theta=17.23, |
||
| 842 | no beam for the first 100k triggers, beam came back 23:47, |
||
| 843 | daq rate is 1.2kHz @ I=25uA |
||
| 844 | |||
| 845 | 3224 production, 00:06-00:30, target=Loop 2 LH2, shms_p=-6,shms_theta=16.17,hms_p=-5.587,hms_theta=17.23, |
||
| 846 | daq rate is 1.280Hz @ I=25uA |
||
| 847 | |||
| 848 | 00:30 target changed to dummy |
||
| 849 | |||
| 850 | 3225 production, 00:32-01:24, target=dummy, shms_p=-6,shms_theta=16.17,hms_p=-5.587,hms_theta=17.23, |
||
| 851 | no beam for the first 100k triggers, beam came back 00:35,not staple until 00:42 |
||
| 852 | daq rate 1.8 kHz @ I=40uA |
||
| 853 | |||
| 854 | 3226 production, 01:24-01:36, target=dummy, shms_p=-6,shms_theta=16.17,hms_p=-5.587,hms_theta=17.23, |
||
| 855 | |||
| 856 | 01:38 SHMS rotated to 19.20deg, HMS rotated to 19.505deg |
||
| 857 | |||
| 858 | 3227 production, 01:39-03:02, target=dummy, shms_p=-5.646,shms_theta=19.28,hms_p=-5.587,hms_theta=19.5, |
||
| 859 | no beam for the first 100k triggers, beam came back 02:42, |
||
| 860 | |||
| 861 | 03:05 changed target to Loop 2 LH2 |
||
| 862 | |||
| 863 | 3228 production, 03:04-03:48, target=Loop 2 LH2,shms_p=-5.646,shms_theta=19.28,hms_p=-5.587,hms_theta=19.5, |
||
| 864 | daq rate 3.7 kHz @ I=73uA |
||
| 865 | daq rate 3.9 kHz @ I=77uA |
||
| 866 | |||
| 867 | 03:51 SHMS Rotated to 23.3 degrees, HMS Rotated to 21.875 degrees |
||
| 868 | |||
| 869 | 3229 production, 03:53-5:44, target=Loop 2 LH2, shms_p=-4.902,shms_theta=23.2,hms_p=-5.587,hms_theta=21.875, |
||
| 870 | no beam for the first 100k triggers, beam came back at 04:37, |
||
| 871 | daq rate 3.9 kHz @ I=77uA |
||
| 872 | |||
| 873 | 05:45 changed target to dummy |
||
| 874 | |||
| 875 | 3230 production, 05:48-06:04, target=dummy, shms_p=-4.902,shms_theta=23.2,hms_p=-5.587,hms_theta=21.875, |
||
| 876 | no beam for the first 100k triggers, beam came back at 05:53, stopped at 5:58 |
||
| 877 | daq rate 1.8 kHz @ I=40uA |
||
| 878 | |||
| 879 | 3231 production, 6:06-06:18, target=dummy, shms_p=-4.902,shms_theta=23.2,hms_p=-5.587,hms_theta=21.875, |
||
| 880 | daq rate 1.8 kHz @ I=40uA |
||
| 881 | |||
| 882 | 06:18 Rotated SHMS to 25.1 deg, rotated HMS to 40 deg, change momenta for both of them, |
||
| 883 | |||
| 884 | 3232 pedestal run, internal pulse, no beam since 06:18 |
||
| 885 | Something went wrong, Hall C configuration change could not finish within 45 minutes and further more |
||
| 886 | the injector has some problem. HallC has been scheduled a controlled access up to six hours from 8AM |
||
| 887 | for the aerogel exchange. |
||
| 888 | |||
| 889 | 11:46 target changed to Loop 2 LH2 |
||
| 890 | 12:40 SHMS rotated to 18.9deg, HMS rotaed to 12.11deg |
||
| 891 | |||
| 892 | 3233 production, 13:36-14:03, target=Loop 2 LH2, shms_p=2.792,shms_theta=18.9,hms_p=-5.587,hms_theta=12.11, |
||
| 893 | daq rate 4.0 kHz @ I=79uA |
||
| 894 | |||
| 895 | 3234 production, 14:03-15:10, target=Loop 2 LH2, shms_p=2.792,shms_theta=18.9,hms_p=-5.587,hms_theta=12.11, |
||
| 896 | daq rate 3.9 kHz @ I=77uA |
||
| 897 | |||
| 898 | 3235 production, 15:11-15:53, target=Loop 2 LH2, shms_p=2.792,shms_theta=18.9,hms_p=-5.587,hms_theta=12.11, |
||
| 899 | |||
| 900 | 3236 production, 15:53-16:15, target=Loop 2 LH2, shms_p=2.792,shms_theta=18.9,hms_p=-5.587,hms_theta=12.11, |
||
| 901 | |||
| 902 | 3237 production, 16:16-17:58, target=Loop 2 LH2, shms_p=2.792,shms_theta=18.9,hms_p=-5.587,hms_theta=12.11, |
||
| 903 | no beam for the first 100k triggers, beam came back at 16:40, |
||
| 904 | daq rate 4.0 kHz @ I=79uA |
||
| 905 | |||
| 906 | 18:00 changed target to dummy |
||
| 907 | |||
| 908 | 3238 production, 18:02-18:16, target=dummy, shms_p=2.792,shms_theta=18.9,hms_p=-5.587,hms_theta=12.11, |
||
| 909 | no beam for the first 100k triggers, beam came back at 18:04, |
||
| 910 | daq rate 1.9 kHz @ I=40uA |
||
| 911 | |||
| 912 | 3239 production, 18:17-19:09, target=dummy, shms_p=2.792,shms_theta=18.9,hms_p=-5.587,hms_theta=12.11, |
||
| 913 | daq rate 1.9 kHz @ I=40uA |
||
| 914 | |||
| 915 | 19:10 target changed to Loop 2 10 cm LH2 |
||
| 916 | 19:12 SHMS rotated to 21.22deg |
||
| 917 | |||
| 918 | 3240 production, 19:14-20:38, target=Loop 2 LH2, shms_p=2.792,shms_theta=21.22,hms_p=-5.587,hms_theta=12.11, |
||
| 919 | daq rate 4.0 kHz @ I=79uA |
||
| 920 | |||
| 921 | 3241 production, 20:39-21:41, target=Loop 2 LH2, shms_p=2.792,shms_theta=21.22,hms_p=-5.587,hms_theta=12.11, |
||
| 922 | daq rate 4.0 kHz @ I=79uA |
||
| 923 | 21:35 target changed to dummy, beam came back at 21:40, |
||
| 924 | the last one minutes is dummy events, I=40uA, daq rate 1.8kHz |
||
| 925 | |||
| 926 | 3242 production, 21:41-22:14, target=dummy, shms_p=2.792,shms_theta=21.22,hms_p=-5.587,hms_theta=12.11, |
||
| 927 | I=40uA, daq rate 1.8kHz |
||
| 928 | |||
| 929 | 22:15 target changed to Loop 2 LH2, SHMS rotated to 23.215deg, actually it is 22.9deg |
||
| 930 | |||
| 931 | 3243 production, 22:17-23:28, target=LH2, shms_p=2.792,shms_theta=22.9,hms_p=-5.587,hms_theta=12.11, |
||
| 932 | I=76uA, daq rate 3.9kHz |
||
| 933 | |||
| 934 | 3244 production, 23:28-00:05, target=LH2, shms_p=2.792,shms_theta=22.9,hms_p=-5.587,hms_theta=12.11, I=80uA |
||
| 935 | 23:35 beam stopped, changed to dummy target, beam came back at 23:44, I=40uA |
||
| 936 | subrun 0 --> LH2 |
||
| 937 | subrun 1 --> part LH2 and part dummy |
||
| 938 | subrun 2 --> dummy |
||
| 939 | |||
| 940 | 00:11 SHMS rotated to 21.22 deg on camera(true angle is 20.9 after subtracting offset), target changed to LH2 |
||
| 941 | The SHMS angle in HallC log is 22.17, that is a typo |
||
| 942 | |||
| 943 | 3245 production, 00:05-00:51, target=LH2, shms_p=2.792,shms_theta=20.9,hms_p=-5.587,hms_theta=12.11 |
||
| 944 | no beam for the first 100k triggers, beam came back at 00:16, |
||
| 945 | beam went away at 00:25 due to accelerator issue, 129k events |
||
| 946 | |||
| 947 | 3246 production, 00:59-04:21, target=LH2, shms_p=2.792,shms_theta=20.9,hms_p=-5.587,hms_theta=12.11 |
||
| 948 | no beam for the first 100k triggers, beam came back at 03:15, |
||
| 949 | 1kHz @ I=20uA |
||
| 950 | 3.2kHz @ I=65uA |
||
| 951 | |||
| 952 | 3247 production, 04:21-05:02, target=LH2, shms_p=2.792,shms_theta=20.9,hms_p=-5.587,hms_theta=12.11 |
||
| 953 | 1.4kHz @ I=29uA |
||
| 954 | |||
| 955 | 05:03 changed to dummy target |
||
| 956 | |||
| 957 | 3248 production, 05:07-05:36, target=dummy, shms_p=2.792,shms_theta=20.9,hms_p=-5.587,hms_theta=12.11 |
||
| 958 | 1.4kHz @ I=30uA |
||
| 959 | |||
| 960 | |||
| 961 | 3249 production, 05:37-10:28, target=dummy, shms_p=2.792,shms_theta=20.9,hms_p=-5.587,hms_theta=12.11 |
||
| 962 | no beam for the first 100k triggers, beam came back at 5:55 last till 6:03. |
||
| 963 | Since 6:03, HMS Q1 was tripped, Hall C could not take beam. |
||
| 964 | 06:37 target changed to LH2 |
||
| 965 | 08:03, beam was back again, I=12uA, target=LH2, shms_p=2.792,shms_theta=20.9,hms_p=-5.587,hms_theta=12.11, |
||
| 966 | 09:10, beam stopped for pass changed |
||
| 967 | subrun 0 has both dummy and LH2 target |
||
| 968 | subrun 1 and 2 are for LH2 target only |
||
| 969 | |||
| 970 | |||
| 971 | ======================================================================================================= |
||
| 972 | 09:15 August 16th, started to change to 3-pass beam , E = 6.399 GeV |
||
| 973 | |||
| 974 | 09:30 SHMS rotated to 17.91 deg, camera shows 18.23 due to the encoder offset 0.32. |
||
| 975 | |||
| 976 | 12:00 Reset SHMS angle offset to zero so the camera shows the same value as the encoder. |
||
| 977 | Set the SHMS to 17.86 degrees and the HMS to 18.31. |
||
| 978 | 3250 pedestal, pulse run, cosmic 12:42 August 16th - 12:05 August 17th, 100203 events |
||
| 979 | 3251 pedestal, pulse run, cosmic, 12:41-14:38, 1000021 events |
||
| 980 | |||
| 981 | memeory upgraded to 24G in daq machine tunl11.jlab.org |
||
| 982 | |||
| 983 | tune beam came back to Hall C |
||
| 984 | 17:17 target changed to carbon hole |
||
| 985 | |||
| 986 | 3252 16:58-17:04 pedestal, pulse run, |
||
| 987 | 3253-3255 debug daq, trash |
||
| 988 | |||
| 989 | 17:48 target changed to Loop2 LH2 |
||
| 990 | |||
| 991 | 3256 17:52-18:01 pedestal, pulse run, debug daq, good for pedestal only |
||
| 992 | |||
| 993 | |||
| 994 | 3257 production, 18:08-18:16, target=LH2,shms_p=2.412,shms_theta=17.86,hms_p=-3.921,hms_theta=13.81, |
||
| 995 | |||
| 996 | 3258 production, 18:22-19:35, target=LH2,shms_p=2.412,shms_theta=17.86,hms_p=-3.921,hms_theta=13.81, |
||
| 997 | daq rate 1152Hz @ I=25uA |
||
| 998 | daq rate 2.3kHz @ I=50uA |
||
| 999 | |||
| 1000 | 19:36 target changed to dummy |
||
| 1001 | |||
| 1002 | 3259 production, 19:36-20:02, target=dummy,shms_p=2.412,shms_theta=17.86,hms_p=-3.921,hms_theta=13.81, |
||
| 1003 | no beam for the first 100k triggers, beam came back at 19:40, |
||
| 1004 | daq rate 1664Hz @ I=40uA |
||
| 1005 | |||
| 1006 | 20:02 SHMS rotated to 19.86 deg |
||
| 1007 | |||
| 1008 | 3260 production, 20:03-20:24, target=dummy, shms_p=2.412,shms_theta=19.86,hms_p=-3.921,hms_theta=13.81, |
||
| 1009 | no beam for the first 100k triggers, beam came back at 20:07, |
||
| 1010 | daq rate 1664Hz @ I=40uA |
||
| 1011 | |||
| 1012 | 20:24 target changed to LH2 |
||
| 1013 | |||
| 1014 | 3261 production, 20:26-20:35, target=LH2, shms_p=2.412,shms_theta=19.86,hms_p=-3.921,hms_theta=13.81, |
||
| 1015 | no beam for the first 100k triggers, beam came back at 20:28, |
||
| 1016 | |||
| 1017 | 3262 production, 20:42-21:01, target=Loop 2 LH2, shms_p=2.412,shms_theta=19.86,hms_p=-3.921,hms_theta=13.81, |
||
| 1018 | daq rate 2.7kHz @ I=60uA |
||
| 1019 | |||
| 1020 | 3263 production, 21:03-21:39, target=Loop 2 LH2, shms_p=2.412,shms_theta=19.86,hms_p=-3.921,hms_theta=13.81, |
||
| 1021 | no beam for the first 100k triggers, beam came back at 21:16, |
||
| 1022 | daq rate 2.7kHz @ I=60uA |
||
| 1023 | |||
| 1024 | 3264 production, 21:39-21:52, target=Loop 2 LH2, shms_p=2.412,shms_theta=19.86,hms_p=-3.921,hms_theta=13.81, |
||
| 1025 | daq rate 2.7kHz @ I=60uA |
||
| 1026 | |||
| 1027 | 3265 production, 21:54-22:23, target=Loop 2 LH2, shms_p=2.412,shms_theta=19.86,hms_p=-3.921,hms_theta=13.81, |
||
| 1028 | daq rate 2.8kHz @ I=60uA |
||
| 1029 | |||
| 1030 | -------- 22:25 SHMS spectrometer change to 21.85------------- |
||
| 1031 | |||
| 1032 | 3266 production, 22:28-22:47, target=Loop 2 LH2, shms_p=2.412,shms_theta=21.85,hms_p=-3.921,hms_theta=13.81, |
||
| 1033 | daq rate 3.2kHz @ I=70uA. Note, it seems they were pushing to reach 80 uA, so the current and rate is may fluctuate |
||
| 1034 | 3267 production, 22:47-23:14, target=Loop 2 LH2, shms_p=2.412,shms_theta=21.85,hms_p=-3.921,hms_theta=13.81 |
||
| 1035 | daq rate 3.6kHz @ I=80uA |
||
| 1036 | no beam for the first 100k triggers, beam came back at 22:49 |
||
| 1037 | 3268 production, 23:15-23:43, target=Loop 2 LH2, shms_p=2.412,shms_theta=21.85,hms_p=-3.921,hms_theta=13.81 |
||
| 1038 | daq rate 3.6kHz @ I=80uA |
||
| 1039 | 3269 production, 23:43-23:55, target=Loop 2 LH2, shms_p=2.412,shms_theta=21.85,hms_p=-3.921,hms_theta=13.81 |
||
| 1040 | daq rate 3.6kHz @ I=80uA |
||
| 1041 | 3270 production, 23:56-24:04, target=Loop 2 LH2, shms_p=2.412,shms_theta=21.85,hms_p=-3.921,hms_theta=13.81 |
||
| 1042 | daq rate 3.60kHz @ I=0uA |
||
| 1043 | no beam for the first 100k triggers, beam was off for 10 minutes due to accelerator issues |
||
| 1044 | |||
| 1045 | 3271 production, 00:07-00:15, target=Loop 2 LH2, shms_p=2.412,shms_theta=21.85,hms_p=-3.921,hms_theta=13.81 |
||
| 1046 | daq rate 3.6kHz @ I=80uA |
||
| 1047 | |||
| 1048 | ---------- 00:20 Switch target to Dummy --------- |
||
| 1049 | |||
| 1050 | 3272 production, 00:19-00:45, target=10 cm Dummy, shms_p=2.412,shms_theta=21.85,hms_p=-3.921,hms_theta=13.81 |
||
| 1051 | no beam for the first 100k triggers, beam returned at 00:24 |
||
| 1052 | daq rate 1.6kHz @ I=40uA |
||
| 1053 | |||
| 1054 | ---------00:50 Switch target to LH2------- |
||
| 1055 | |||
| 1056 | 3273 production, 00:56-02:22, target=Loop 2 LH2, shms_p=2.412,shms_theta=21.85,hms_p=-3.921,hms_theta=13.8 |
||
| 1057 | daq rate 1.8kHz @ I=40uA |
||
| 1058 | |||
| 1059 | |||
| 1060 | 3274 production, 02:23-3:04, target=Loop 2 LH2, shms_p=2.412,shms_theta=21.86,hms_p=-3.921,hms_theta=13.8 |
||
| 1061 | daq rate 850Hz @ I=20uA |
||
| 1062 | daq rate 512Hz @ I=12uA |
||
| 1063 | |||
| 1064 | 3275 production, 03:04-04:44, target=Loop 2 LH2, shms_p=2.412,shms_theta=21.86,hms_p=-3.921,hms_theta=13.8 |
||
| 1065 | daq rate 512Hz @ I=12uA |
||
| 1066 | |||
| 1067 | |||
| 1068 | 3276 production, 04:44-8:49, target=Loop 2 LH2, shms_p=2.412,shms_theta=21.86,hms_p=-3.921,hms_theta=13.8 |
||
| 1069 | daq rate 512Hz @ I=12uA |
||
| 1070 | HallC was doing FADC dead time study. current changed from 12uA, to 60uA then to 30uA |
||
| 1071 | |||
| 1072 | 06:27 HMS rotated to 18.595deg, SHMS roated to 37.97deg |
||
| 1073 | 07:00 I=75UA, target=LH2, shms_p=2.412,shms_theta=38,hms_p=-4.752,hms_theta=18.61, |
||
| 1074 | 07:28 target changed to dummy, I=40uA |
||
| 1075 | 07:48 HMS rotated to 21.095deg, SHMS rotated to 34.465deg |
||
| 1076 | target=dummy, shms_p=2.792,shms_theta=34.49,hms_p=-4.391,hms_theta=21.11, |
||
| 1077 | 08:15 target changed to LH2, I=75uA |
||
| 1078 | target=LH2, shms_p=2.792,shms_theta=34.49,hms_p=-4.391,hms_theta=21.11, |
||
| 1079 | |||
| 1080 | subrun 0-6: SHMS=21.86deg |
||
| 1081 | subrun 7: SHMS=21.86deg and 37.97deg |
||
| 1082 | subrun 8-11: SHMS=37.97deg |
||
| 1083 | subrun 12: SHMS=37.97deg and 34.49deg |
||
| 1084 | subrun 13-16: SHMS=34.49deg |
||
| 1085 | |||
| 1086 | |||
| 1087 | 08:59 HMS rotated to 26.55deg, SHMS rotated to 28.46deg |
||
| 1088 | |||
| 1089 | 3277 production, 08:53-10:45, target=LH2, shms_p=3.493,shms_theta=28.47,hms_p=-3.727,hms_theta=26.56, |
||
| 1090 | no beam for the first 100k triggers, beam returned at 09:30, I=75uA |
||
| 1091 | 10:10 target changed to dummy, I=40uA |
||
| 1092 | 10:27 beam stopped |
||
| 1093 | |||
| 1094 | ===================20020818 Install VTP board========================= |
||
| 1095 | From here EPICS variables are inserted into ER |
||
| 1096 | |||
| 1097 | SHMS at 7.66deg and HMS at 28.78deg. |
||
| 1098 | |||
| 1099 | 3278-3283 debug daq runs. HV are off. no beam. good for pedestal |
||
| 1100 | HV are turned on in the middle of run 3283, but no beam after HV is on. |
||
| 1101 | |||
| 1102 | 3284 production, 20:20-23:14, target=LH2,shms_p=4.605,shms_theta=7.66,hms_p=-1.727,hms_theta=28.78, |
||
| 1103 | no beam for the first 100k triggers, beam came back at 21:11 |
||
| 1104 | daq rate 750Hz @ I=75UA |
||
| 1105 | 22:00 target changed to dummy |
||
| 1106 | 22:40 SHMS rotated to 9.66deg, target changed to LH2 |
||
| 1107 | target=LH2,shms_p=4.605,shms_theta=9.66,hms_p=-1.727,hms_theta=28.78, |
||
| 1108 | daq rate 1235Hz @ I=75UA |
||
| 1109 | subrun 0 contains data for SHMS=7.66deg |
||
| 1110 | subrun 1 contains data for both 7.66deg and 9.66 |
||
| 1111 | subrun 2-3 contains data for SHMS=9.66deg |
||
| 1112 | |||
| 1113 | 3285 production, 23:19-00:52 August 19th, target=LH2,shms_p=4.605,shms_theta=9.66,hms_p=-1.727,hms_theta=28.78, |
||
| 1114 | daq rate 1235Hz @ I=75UA |
||
| 1115 | |||
| 1116 | 3286 production, 00:54-05:28, target=LH2,shms_p=4.605,shms_theta=9.66,hms_p=-1.727,hms_theta=28.78, |
||
| 1117 | coda lost roc3 at 01:15 and did not take data since then. |
||
| 1118 | 03:00 changed target to dummy |
||
| 1119 | 03:23 SHMS rotated to 11.66deg, |
||
| 1120 | 03:50 target changed to LH2,shms_p=4.605,shms_theta=11.67,hms_p=-1.727,hms_theta=28.78, |
||
| 1121 | |||
| 1122 | |||
| 1123 | 3287 production, 05:29-12:57,target=LH2, shms_p=4.605,shms_theta=11.67,hms_p=-1.727,hms_theta=28.78, |
||
| 1124 | 8:25 SHMS rotated to 10deg and HMS to 42.23deg |
||
| 1125 | target=LH2,shms_p=5.127,shms_theta=10,hms_p=-1.158,hms_theta=42.26, |
||
| 1126 | no beam between 10:09 and 11:20 |
||
| 1127 | daq rate 1.5kHz @ I=77uA |
||
| 1128 | subrun 0-19: 11.67deg |
||
| 1129 | subrun 20: 11.67deg and 10deg |
||
| 1130 | subrun 21-29: 10deg |
||
| 1131 | |||
| 1132 | |||
| 1133 | 3288 production, 12:57-18:12, target=LH2, shms_p=5.127,shms_theta=10,hms_p=-1.158,hms_theta=42.26, |
||
| 1134 | daq rate 1.5kHz @ I=77uA |
||
| 1135 | |||
| 1136 | 3289 production, 18:14-18:31, target=LH2, shms_p=5.127,shms_theta=10,hms_p=-1.158,hms_theta=42.26, |
||
| 1137 | daq rate 1.5kHz @ I=77uA |
||
| 1138 | |||
| 1139 | 3290 production, 18:31-00:06 August 20th, target=LH2, shms_p=5.127,shms_theta=10,hms_p=-1.158,hms_theta=42.26, |
||
| 1140 | daq rate 1.5kHz @ I=77uA |
||
| 1141 | 22:09 target changed to dummy, I=40uA |
||
| 1142 | daq rate 512Hz @ I=40uA |
||
| 1143 | |||
| 1144 | 3291 production, 00:08-00:34 August 20th, target=dummy, shms_p=5.127,shms_theta=10,hms_p=-1.158,hms_theta=42.26, |
||
| 1145 | daq rate 512Hz @ I=40uA |
||
| 1146 | 0:26 target changed to LD2 |
||
| 1147 | Ion chamber calibration start from 00:31 |
||
| 1148 | |||
| 1149 | 3292 production, 00:34-00:49 August 20th, target=LD2, shms_p=5.127,shms_theta=10,hms_p=-1.158,hms_theta=42.26, |
||
| 1150 | beam current ramped up from 0 to 32uA for Ion chamber calibration |
||
| 1151 | |||
| 1152 | 3293 production, 00:50-08:50 August 20th, target=LD2, shms_p=5.127,shms_theta=10,hms_p=-1.158,hms_theta=42.26, |
||
| 1153 | daq rate 3.2kHz @ I=77uA |
||
| 1154 | |||
| 1155 | 3294 production, 08:51-14:18 August 20th, target=LD2, shms_p=5.127,shms_theta=10,hms_p=-1.158,hms_theta=42.26, |
||
| 1156 | daq rate 3.2kHz @ I=77uA |
||
| 1157 | 10:36 SHMS rotated to 8deg, |
||
| 1158 | target=LD2, shms_p=5.127,shms_theta=8,hms_p=-1.158,hms_theta=42.26, daq rate 2.0kHz @ I=78.5uA |
||
| 1159 | |||
| 1160 | subrun 0-12: SHMS=10 deg |
||
| 1161 | subrun 13: SHMS=10 deg and 8 deg |
||
| 1162 | subrun 14-30: SHMS=8deg |
||
| 1163 | |||
| 1164 | 3295 production, 14:18-01:19 August 21st, target=LD2, shms_p=5.127,shms_theta=8,hms_p=-1.158,hms_theta=42.26, |
||
| 1165 | daq rate 2.0kHz @ I=78.5uA |
||
| 1166 | 00:18 taret changed to dummy, I=40uA |
||
| 1167 | |||
| 1168 | subrun 0-48: LD2 |
||
| 1169 | subrun 49: LD2 and dummy |
||
| 1170 | subrun 50: dummy |
||
| 1171 | |||
| 1172 | 3296 production, 14:18-01:21 August 21st, target=dummy, shms_p=5.127,shms_theta=8,hms_p=-1.158,hms_theta=42.26, |
||
| 1173 | just 106k events |
||
| 1174 | |||
| 1175 | 3297 production, 01:24-02:54 August 21st, target=dummy, shms_p=5.127,shms_theta=8,hms_p=-1.158,hms_theta=42.26, |
||
| 1176 | daq rate 500Hz @ I=40uA |
||
| 1177 | 02:36 changed to LH2 target, I=80uA |
||
| 1178 | |||
| 1179 | |||
| 1180 | 3298 production, 02:55-14:46 August 21st, target=LH2, shms_p=5.127,shms_theta=8,hms_p=-1.158,hms_theta=42.26, |
||
| 1181 | daq rate 1000Hz @ I=78uA |
||
| 1182 | |||
| 1183 | 3299 production, 14:46-19:00 August 21st, target=LH2, shms_p=5.127,shms_theta=8,hms_p=-1.158,hms_theta=42.26, |
||
| 1184 | daq rate 1000Hz @ I=78uA |
||
| 1185 | |||
| 1186 | 19:01 SHMS rotated to 10deg |
||
| 1187 | |||
| 1188 | 3300 production, 19:02-00:57 August 22th, target=LH2, shms_p=5.127,shms_theta=10,hms_p=-1.158,hms_theta=42.26, |
||
| 1189 | daq rate 1500Hz @ I=78uA |
||
| 1190 | 23:47 target changed to LD2 |
||
| 1191 | |||
| 1192 | 3301 production, 00:57-02:51 August 22nd, target=LD2, shms_p=5.127,shms_theta=10,hms_p=-1.158,hms_theta=42.26, |
||
| 1193 | daq rate 3.5 kHz @ I=80uA |
||
| 1194 | beam stopped at 02:41 |
||
| 1195 | |||
| 1196 | 02:55 target changed to LH2 |
||
| 1197 | 03:00 SHMS rotated to 7.5deg, HMS rotated to 29.55deg |
||
| 1198 | |||
| 1199 | 03:00 HMS Q1 interlock happened and SHMS Dipole did not response. Hall is in controlled access. |
||
| 1200 | 05:00 Accelerator fire alarm, no beam. |
||
| 1201 | |||
| 1202 | 3302 pedestal, 02:51-09:09 August 22nd, no beam |
||
| 1203 | |||
| 1204 | 10:50 SHMS rotated to 5.5deg, |
||
| 1205 | |||
| 1206 | 3303 production, 15:01-01:04 August 23rd, target=LH2, shms_p=5.422,shms_theta=5.49,hms_p=-0.96,hms_theta=29.57, |
||
| 1207 | no beam until 16:20 |
||
| 1208 | daq rate 500Hz @ I=40uA |
||
| 1209 | |||
| 1210 | 3304 production, 01:04-04:02 August 23rd, target=LH2, shms_p=5.422,shms_theta=5.49,hms_p=-0.96,hms_theta=29.57, |
||
| 1211 | daq rate 500Hz @ I=40uA |
||
| 1212 | 03:00 changed to dummy target |
||
| 1213 | beam stop at 03:15, Trim rack W301B13 negative bulk supply fail |
||
| 1214 | |||
| 1215 | 3305 production, 04:04-11:33 August 23rd, target=dummy, shms_p=5.422,shms_theta=5.49,hms_p=-0.96,hms_theta=29.57, |
||
| 1216 | no beam at the beginning until 04:58 |
||
| 1217 | 06:25 target changed to LD2 |
||
| 1218 | 11:23 target changed to 'home' (absolute empty) for BCM calibration |
||
| 1219 | |||
| 1220 | |||
| 1221 | ================================================================================== |
||
| 1222 | run 3171-3305 were using the same HV as FermiLab, but the HV for Shower_L is too low, |
||
| 1223 | the amplitude of the signal is about 10mv. |
||
| 1224 | |||
| 1225 | Here is the new HVs: |
||
| 1226 | Slot 2 |
||
| 1227 | Name: Shower_L(SDU#4), Shower_R(THU#2), Shower_T(SDU#5), CherA, CherB, CherC, CherD |
||
| 1228 | HV: 1080 1100 930 1050 1050 1050 1050 |
||
| 1229 | |||
| 1230 | Slot 10 |
||
| 1231 | Name: SC0_T, SC0_B, SC1_T, SC1_B, LASPD_T, LASPD_B, SPD, Preshower_L(SDU#1), Preshower_R(THU#1), Preshower_T(SDU#2) |
||
| 1232 | HV: 1600 1600 1700 1750 850 1000 1950 1550 1800 1750 |
||
| 1233 | |||
| 1234 | |||
| 1235 | 12:54 target changed to Loop 1 10cm LD2 |
||
| 1236 | |||
| 1237 | 3306 procudtion, 12:55-14:20 August 23rd, target=LD2, shms_p=5.422,shms_theta=5.49,hms_p=-0.96,hms_theta=29.57, |
||
| 1238 | no beam for the first 100k triggers, beam came back at 13:18 |
||
| 1239 | daq rate = 2.7kHz @ I=31uA |
||
| 1240 | daq rate = 3.0kHz @ I=33uA |
||
| 1241 | This daq rate is too high comparing to run 3307. The beam might scratch the target frame. |
||
| 1242 | |||
| 1243 | ==========HV changed======= |
||
| 1244 | Lower HVs for preshower_top and all 3 Ecal blocks |
||
| 1245 | |||
| 1246 | Slot 2 |
||
| 1247 | Name: Shower_L(SDU#4), Shower_R(THU#2), Shower_T(SDU#5), CherA, CherB, CherC, CherD |
||
| 1248 | HV: 1030 1050 900 1050 1050 1050 1050 |
||
| 1249 | |||
| 1250 | Slot 10 |
||
| 1251 | Name: SC0_T, SC0_B, SC1_T, SC1_B, LASPD_T, LASPD_B, SPD, Preshower_L(SDU#1), Preshower_R(THU#1), Preshower_T(SDU#2) |
||
| 1252 | HV: 1600 1600 1700 1750 850 1000 1950 1550 1800 1700 |
||
| 1253 | |||
| 1254 | |||
| 1255 | 3307 procudtion, 14:23-18:23 August 23rd, target=LD2, shms_p=5.422,shms_theta=5.49,hms_p=-0.96,hms_theta=29.57, |
||
| 1256 | daq rate = 1.7kHz @ I=31uA |
||
| 1257 | |||
| 1258 | |||
| 1259 | ==========HV changed======= |
||
| 1260 | 18:24 Adjusted TOP preshower (-40V) and top shower HVs (-30V) |
||
| 1261 | |||
| 1262 | Slot 2 |
||
| 1263 | Name: Shower_L(SDU#4), Shower_R(THU#2), Shower_T(SDU#5), CherA, CherB, CherC, CherD |
||
| 1264 | HV: 1030 1050 870 1050 1050 1050 1050 |
||
| 1265 | |||
| 1266 | Slot 10 |
||
| 1267 | Name: SC0_T, SC0_B, SC1_T, SC1_B, LASPD_T, LASPD_B, SPD, Preshower_L(SDU#1), Preshower_R(THU#1), Preshower_T(SDU#2) |
||
| 1268 | HV: 1600 1600 1700 1750 850 1000 1950 1550 1800 1660 |
||
| 1269 | |||
| 1270 | |||
| 1271 | 3308 procudtion, 18:25-18:34 August 23rd, target=LD2, shms_p=5.422,shms_theta=5.49,hms_p=-0.96,hms_theta=29.57, |
||
| 1272 | daq rate = 1.0kHz @ I=31uA |
||
| 1273 | |||
| 1274 | ==========HV changed======= |
||
| 1275 | SC1_top 1700 -> 1680 |
||
| 1276 | Top preshower 1660-> 1640 |
||
| 1277 | Top SHower 870 -> 850 |
||
| 1278 | |||
| 1279 | 06:45 SHMS rotated to 7.5 deg |
||
| 1280 | |||
| 1281 | 3309 procudtion, 18:48-20:08 August 23rd, target=LD2, shms_p=5.422,shms_theta=7.5,hms_p=-0.96,hms_theta=29.57, |
||
| 1282 | no beam for the first 100k triggers, beam came back at 19:36 |
||
| 1283 | beam stopped at 19:54 due to accelerator issue. |
||
| 1284 | |||
| 1285 | 3310 production, 20:11-03:17 August 24th, target=LD2, shms_p=5.422,shms_theta=7.5,hms_p=-0.96,hms_theta=29.57, |
||
| 1286 | daq rate 1.7kHz @I=80uA |
||
| 1287 | |||
| 1288 | 3311 procudtion, 03:17-12:33 August 24th, target=LD2, shms_p=5.422,shms_theta=7.5,hms_p=-0.96,hms_theta=29.57, |
||
| 1289 | no beam for the first 100k triggers, beam came back at 03:18 |
||
| 1290 | daq rate 1.7kHz @I=80uA |
||
| 1291 | 8:38 changed target to dummy, I=40uA |
||
| 1292 | 10:40 changed target to 'home'(absolute empty) for beam energy measurement, I=100nA, daq rate=1.2kHz |
||
| 1293 | |||
| 1294 | 3312 trash, 12:38-12:44 August 24th, target=empty(?), shms_p=5.422,shms_theta=7.5,hms_p=-0.96,hms_theta=29.57, |
||
| 1295 | daq rate=1.2kHz @ I=0.1uA, the beam must have scratched something, 173k events |
||
| 1296 | beam energy measurement result: beam = 6394.694 +/- 3.055 MeV |
||
| 1297 | |||
| 1298 | 15:42 changed target to Loop2 LH2 |
||
| 1299 | |||
| 1300 | 3313 procudtion, 12:52-01:06 August 25th, target=LH2, shms_p=5.422,shms_theta=7.5,hms_p=-0.96,hms_theta=29.57, |
||
| 1301 | no beam for the first 100k triggers, beam came back at 15:38 |
||
| 1302 | daq rate=750Hz @ I=78uA, |
||
| 1303 | |||
| 1304 | 3314 procudtion, 01:07-08:56 August 25th, target=LH2, shms_p=5.422,shms_theta=7.5,hms_p=-0.96,hms_theta=29.57, |
||
| 1305 | daq rate=750Hz @ I=78uA, |
||
| 1306 | 03:00 Rotated SHMS to 26.82 deg, HMS to 28.41 deg, shms_p=3.706,shms_theta=26.82,hms_p=-3.514,hms_theta=28.41, |
||
| 1307 | 04:16 changed target to dummy, I=40uA |
||
| 1308 | 04:35 Rotated SHMS to 23.13 deg, HMS to 33.37 deg, shms_p=4.22,shms_theta=23.13,hms_p=-3.014,hms_theta=33.37, |
||
| 1309 | 05:15 changed target to LH2 |
||
| 1310 | |||
| 1311 | |||
| 1312 | 3315 procudtion, 08:57-10:27 August 25th, target=LH2, shms_p=4.22,shms_theta=23.13,hms_p=-3.014,hms_theta=33.37, |
||
| 1313 | daq rate=3.0kHz @ I=78uA, |
||
| 1314 | beam stopped at 09:03, 777k evnets |
||
| 1315 | |||
| 1316 | 09:30 Rotated SHMS to 8.0 deg and HMS to 42.23deg |
||
| 1317 | |||
| 1318 | SBSVME22 lost connection. USB stick failed. No access until next accelerator maintenance day. |
||
| 1319 | can not control MPOD utill the 20220916. |
||
| 1320 | ============================================================================================== |
||
| 1321 | |||
| 1322 | 3318 pedestal, test daq, no beam, HV on |
||
| 1323 | 3319 pedestal, test daq, no beam, HV on |
||
| 1324 | |||
| 1325 | 20220917 23:23 |
||
| 1326 | |||
| 1327 | 3320 gem pedestal run, GEM HV was not on for this run |
||
| 1328 | |||
| 1329 | 3321 trash run, pulser event enabled, unable to check GEM signal |
||
| 1330 | |||
| 1331 | ==first 100k pulse event are disabled== |
||
| 1332 | |||
| 1333 | 3322 beam seems stopped |
||
| 1334 | 23:51 still no beam |
||
| 1335 | 00:28 no beam |
||
| 1336 | 01:50 still no beam |
||
| 1337 | 02:22 beam delivery suspended due to transport ion chamber failure |
||
| 1338 | |||
| 1339 | 3323 Latency check, initial latency value 49 |
||
| 1340 | 3324 initial latency value 49 -- (49 is too late) |
||
| 1341 | 3325 & 3326 scan for value 52 |
||
| 1342 | 3327 & 3328 scan for value 55 -- (this seems good, also consistent with Ezekiel's INFN setup) |
||
| 1343 | 3329 & 3330 & 3331 & 3332 scan for value 58 (seems too early) |
||
| 1344 | 3333 & 3334 scan for value 64 -- (64 is too early) |
||
| 1345 | 3335 & 3336 scan for value 60 -- (too early) |
||
| 1346 | |||
| 1347 | 3337 production, 00:25 - 11:06 Sep. 19th, run with new GEM latency 55, (55 is too early) |
||
| 1348 | |||
| 1349 | 3338 procudtion, 11:24 Sep 19th - 18:34 Sep. 20th, run with slightly modifed latency 52. |
||
| 1350 | target=Fe, shms_p=-8.55,shms_theta=8.3,hms_p=1.82,hms_theta=48.3, |
||
| 1351 | 13:10 Sep 19th, changed target to B10 |
||
| 1352 | 14:16 Sep 19th, changed target to B4C-11 |
||
| 1353 | 16:00 Sep 19th, beam stopped till 00:24 Sep 20th |
||
| 1354 | 05:03 Sep 20th, changed target to Ca40, HMS roated to 66.4deg, HMS_p was set to 1.325 |
||
| 1355 | target=Ca40,shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, |
||
| 1356 | 09:00 Sep 20th, RF study started, no beam till 21:12 |
||
| 1357 | |||
| 1358 | 3339 production, 01:58 - 12:43 Sep 21st, |
||
| 1359 | beam was bad, accelarator worked at about 50% efficiency |
||
| 1360 | target=Ca40,shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, |
||
| 1361 | 10:44, SHMS Q1 vacuum issiu. Beam stopped. |
||
| 1362 | 11:54, Beam came back but last for 10 minutes only then went away. |
||
| 1363 | |||
| 1364 | 3340 production,13:07 Sep 21st - 12:03 Sep 22nd, |
||
| 1365 | target=Ca40,shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, |
||
| 1366 | |||
| 1367 | 3341 production,12:04 Sep 22nd - 23:35 Sep 22nd, |
||
| 1368 | target=Ca40,shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, |
||
| 1369 | |||
| 1370 | 3342 production,23:36 Sep 22nd - 15:21 Sep 23rd, |
||
| 1371 | target=Ca40,shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, |
||
| 1372 | beam stopped due to vacuum issue at around 9:16. |
||
| 1373 | |||
| 1374 | 3343 production,20:06 Sep 23rd - |
||
| 1375 | target=Ca40,shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, |
||
| 1376 | daq crashed at 02:22 Sep 25th |
||
| 1377 | |||
| 1378 | Tunl11 lost mounting disk, vncserver can not start properly |
||
| 1379 | 3344 daq crashed, no data file |
||
| 1380 | |||
| 1381 | 3345 productin, 21:02 Sep 25th - 03:51 Sep 26th |
||
| 1382 | target=carbon, shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, |
||
| 1383 | daq rate 250Hz |
||
| 1384 | |||
| 1385 | target changed to B4C-10 |
||
| 1386 | 3346 production, 03:59 - 11:18 Sep 26th |
||
| 1387 | target=B4C-10, shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, |
||
| 1388 | beam stopped at ~05:45 |
||
| 1389 | |||
| 1390 | 3347 production, 11:20 Sep 26th - 20:34 Sep 27th, no beam until 20:12 Sep 26th. |
||
| 1391 | target=B4C-10, shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, |
||
| 1392 | 20:30 target changed to Ca48 |
||
| 1393 | target=Ca48,shms_p=-8.55,shms_theta=8.3,hms_p=1.82,hms_theta=48.33, |
||
| 1394 | 03:51 Sep 27th, target changed to Ca40 |
||
| 1395 | 04:19 Sep 27th, target changed to Carbon |
||
| 1396 | 04:35 Sep 27th, target changed to Beryllium |
||
| 1397 | 05:12 Sep 27th, target changed to B4C-10 |
||
| 1398 | 05:40 Sep 27th, target changed to B4C-11 |
||
| 1399 | 06:00 Sep 27th, target changed to Beryllium |
||
| 1400 | 13:42 Sep 27th, target changed to B4C-10 |
||
| 1401 | beam stopped at 17:20, Sep 27th |
||
| 1402 | |||
| 1403 | 3348 production, 20:51, Sep 27th - 15:27 Sep 28th, no beam until 02:49 Sep 28th. |
||
| 1404 | 22:18 Sep 27th, target changed to B4C-11 |
||
| 1405 | target=B4C-11, shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, |
||
| 1406 | 09:59 Sep 28th, target changed to B4C-10 |
||
| 1407 | 14:35 Sep 28th, target changed to Beryllium |
||
| 1408 | |||
| 1409 | 3349 production, 15:28, Sep 28th - 10:40 Sep 29th, no beam until 20:12 Sep 26th. |
||
| 1410 | target=Beryllium, shms_p=-8.55,shms_theta=8.3,hms_p=1.325,hms_theta=66.4, rate=500Hz |
||
| 1411 | 18:51 target changed to Loop 1 LD2 |
||
| 1412 | 20:20 target changed to Fe |
||
| 1413 | beam stopped at 7:00 Sep 29th. |
||
| 1414 | |||
| 1415 | ==TS#3 enabled, which is ShowerSum== |
||
| 1416 | |||
| 1417 | 3350 trash, daq test |
||
| 1418 | |||
| 1419 | |||
| 1420 | 18:14 SHMS rotated to 26deg |
||
| 1421 | |||
| 1422 | 3351 pedestal, 18:21 - 18:27 Sep 29th, no beam |
||
| 1423 | |||
| 1424 | 3352 production, 18:28 Sep 29th - 22:32 Oct 1st, no beam until 22:34 |
||
| 1425 | 23:04 Sep 29th, target changed to Carbon hole |
||
| 1426 | target=Carbon Hole, shms_p=-4.935,shms_theta=26.02,hms_p=6.476,hms_theta=20.69, |
||
| 1427 | 23:30 Sep 29th, beam stopped |
||
| 1428 | 23:31 Sep 29th, target changed to Loop 2 LH2 |
||
| 1429 | 10:04 Oct 1st, tune beam came back |
||
| 1430 | target=LH2, shms_p=-4.935,shms_theta=26.02,hms_p=6.476,hms_theta=20.69, |
||
| 1431 | 15:38 Oct 1st, target changed to Caron |
||
| 1432 | 15:57 Oct 1st, Target changed to LH2 |
||
| 1433 | 17:01 Oct 1st, SHMS at 9.24 deg and HMS at 29.39 deg |
||
| 1434 | target=LH2, shms_p=-8,shms_theta=9.24,hms_p=4.17,hms_theta=29.41, |
||
| 1435 | daq crashed at 18:34 |
||
| 1436 | 20:28 Oct 1st, SHMS rotated to 10.77, HMS to 31.73 |
||
| 1437 | shms_p=-8,shms_theta=10.77,hms_p=3.76,hms_theta=31.73, |
||
| 1438 | |||
| 1439 | subrun 00-01: SHMS=26, Carbon hole + LH2 |
||
| 1440 | subrun 02-23: SHMS=26deg, LH2 |
||
| 1441 | subrun 24: SHMS=26deg, LH2 + carbon target |
||
| 1442 | subrun 25: SHMS=26deg, Carbon |
||
| 1443 | subrun 26: SHMS=26deg, carbon + SHMS=9.24deg LH2 target |
||
| 1444 | subrun 27-42: shms_theta=9.24, LH2 |
||
| 1445 | |||
| 1446 | |||
| 1447 | 3353 production, 22:35 Oct 1st - 23:01 Oct 1st, daq crashed at about 22:55 |
||
| 1448 | target=LH2, shms_p=-8,shms_theta=10.77,hms_p=3.76,hms_theta=31.73, |
||
| 1449 | 22:38 Oct 1st, SHMS rotated to 16.55, HMS to 29.4 |
||
| 1450 | target=LH2, shms_p=-8,shms_theta=16.55,hms_p=4.17,hms_theta=29.4, daq rate=7kHz |
||
| 1451 | |||
| 1452 | 3354 production, 23:02 Oct 1st - 23:40 OCt 1st, |
||
| 1453 | target=LH2, shms_p=-8,shms_theta=16.55,hms_p=4.17,hms_theta=29.4, daq rate=7kHz |
||
| 1454 | 23:25 target changed to dummy |
||
| 1455 | target=dummy, shms_p=-8,shms_theta=16.55,hms_p=4.17,hms_theta=29.4, daq rate=6.7kHz |
||
| 1456 | |||
| 1457 | ==TS#3 prescale factor set to 10== |
||
| 1458 | |||
| 1459 | 3355 production, 23:42 Oct 1st - 20:43 Oct 2nd |
||
| 1460 | target=dummy, shms_p=-8,shms_theta=16.55,hms_p=4.17,hms_theta=29.4, daq rate=2.4kHz |
||
| 1461 | 23:55 Oct 1st, SHMS rotated to 15.09deg |
||
| 1462 | 00:15 Oct 2nd, target change to LH2 |
||
| 1463 | target=LH2,shms_p=-8,shms_theta=15.09,hms_p=3.77,hms_theta=31.72, |
||
| 1464 | 00:51 Oct 2nd, SHMS rotated to 13.67deg |
||
| 1465 | 01:07 Oct 2nd, target change to dummy |
||
| 1466 | target=dummy,shms_p=-8,shms_theta=13.67,hms_p=3.35,hms_theta=34.36, |
||
| 1467 | 01:37 Oct 2nd, SHMS rotated to 12.24deg |
||
| 1468 | target=dummy,shms_p=-8,shms_theta=12.24,hms_p=2.94,hms_theta=37.37, |
||
| 1469 | 02:08 Oct 2nd,target changed to LH2 |
||
| 1470 | 02:25 Oct 2nd, SHMS rotated to 10.78deg |
||
| 1471 | target=LH2,shms_p=-8,shms_theta=10.78,hms_p=2.51,hms_theta=40.97, |
||
| 1472 | 02:57 Oct 2nd,target changed to dummy |
||
| 1473 | 03:25 Oct 2nd, SHMS rotated to 9.25deg |
||
| 1474 | target=dummy,shms_p=-8,shms_theta=9.25,hms_p=2.08,hms_theta=45.36, |
||
| 1475 | 03:42 Oct 2nd,target changed to LH2 |
||
| 1476 | 04:30 Oct 2nd, SHMS rotated to 26.02deg |
||
| 1477 | target=LH2,shms_p=-4.935,shms_theta=26.02,hms_p=6.476,hms_theta=19.55, |
||
| 1478 | 12:29 Oct 2nd,target changed to dummy |
||
| 1479 | 13:30 Oct 2nd, SHMS rotated to 20.02deg |
||
| 1480 | target=dummy,shms_p=-6.286,shms_theta=20.02,hms_p=5.107,hms_theta=24.91, |
||
| 1481 | 14:02 Oct 2nd,target changed to LH2 |
||
| 1482 | 15:30 Oct 2nd, SHMS rotated to 15.01deg |
||
| 1483 | target=LH2,shms_p=-7.626,shms_theta=15.01,hms_p=3.738,hms_theta=31.89, |
||
| 1484 | 16:05 Oct 2nd,target changed to dummy |
||
| 1485 | 16:30 Oct 2nd, SHMS rotated to 10.02deg |
||
| 1486 | target=dummy,shms_p=-9.005,shms_theta=10.02,hms_p=2.289,hms_theta=43.13, |
||
| 1487 | 17:55 Oct 2nd,target changed to LH2 |
||
| 1488 | 18:40 Oct 2nd, SHMS rotated to 10.02deg |
||
| 1489 | target=LH2,shms_p=-9.2,shms_theta=9.26,hms_p=2.076,hms_theta=45.39, |
||
| 1490 | 19:05 Oct 2nd,target changed to dummy |
||
| 1491 | 19:35 Oct 2nd, SHMS rotated to 7.995deg |
||
| 1492 | target=dummy,shms_p=-9.502,shms_theta=7.995,hms_p=1.739,hms_theta=49.53, |
||
| 1493 | 21:30 Oct 2nd,target changed to LH2 |
||
| 1494 | 21:54 Oct 2nd,SJMS sieve in |
||
| 1495 | 22:18 Oct 2nd,target changed to Beryllium, |
||
| 1496 | target=Be,SHMS Sieve Slit in,shms_p=-9.502,shms_theta=7.995,hms_p=1.739,hms_theta=49.53, |
||
| 1497 | |||
| 1498 | |||
| 1499 | 3356 production, 20:44 Oct 2nd - 17:05 Oct 3rd |
||
| 1500 | target=dummy, shms_p=-9.502,shms_theta=7.995,hms_p=1.739,hms_theta=49.53, |
||
| 1501 | 22:52 Oct 2nd,target changed to LH2 |
||
| 1502 | target=LH2,SHMS Sieve Slit in,shms_p=-9.502,shms_theta=7.995,hms_p=1.739,hms_theta=49.53, |
||
| 1503 | 23:28 Oct 2nd,target changed to dummy, SHMS sieve still in |
||
| 1504 | target=dummy,SHMS Sieve Slit in,shms_p=-9.502,shms_theta=7.995,hms_p=1.739,hms_theta=49.53, |
||
| 1505 | 23:45 Oct 2nd, SHMS momentum set to -8.55 |
||
| 1506 | 00:08 Oct 3rd,target changed to LH2 |
||
| 1507 | target=LH2,SHMS Sieve Slit in,shms_p=-8.55,shms_theta=7.995,hms_p=1.739,hms_theta=49.53, |
||
| 1508 | |||
| 1509 | 00:49 Oct 3rd,target changed to Beryllium, |
||
| 1510 | target=Beryllium,SHMS Sieve Slit in,shms_p=-8.55,shms_theta=7.995,hms_p=1.739,hms_theta=49.53, |
||
| 1511 | 01:28 Oct 3rd,target changed to LH2 |
||
| 1512 | 01:42 roated SHMS and HMS. Spectrometer Emergency Stop Detected, Hall is open for trouble shooting. |
||
| 1513 | beam stopped from 1:38. |
||
| 1514 | Accelerator can not deliver beam since 04:10 Oct 3rd |
||
| 1515 | |||
| 1516 | |||
| 1517 | |||
| 1518 | ==SHMS roated to 20.02deg, HMS to 20deg== |
||
| 1519 | |||
| 1520 | 3357 production, 18:20 Oct 3rd - 12:14 Oct 4th |
||
| 1521 | 23:45 Oct 3rd, tune beam |
||
| 1522 | target=LH2,shms_p=-4,shms_theta=20.02,hms_p=-4,hms_theta=20.0, |
||
| 1523 | 01:05 SHMS and HMS rotated to 25deg |
||
| 1524 | target=LH2,shms_p=-3.5,shms_theta=25.02,hms_p=-3.5,hms_theta=25.0, |
||
| 1525 | 03:22 SHMS and HMS rotated to 25deg |
||
| 1526 | target=LH2,shms_p=-2.5,shms_theta=35.02,hms_p=-2.5,hms_theta=35.0, |
||
| 1527 | 05:27 change momentum to 1.24GeV SHMS |
||
| 1528 | target=LH2,shms_p=-1.24,shms_theta=35.02,hms_p=-2.5,hms_theta=35.0, |
||
| 1529 | 08:19 SHMS rotated to 8 deg, momentum is set to -9.2GeV |
||
| 1530 | 08:45 target changed to LD2 |
||
| 1531 | target=LD2,shms_p=-9.2,shms_theta=8,hms_p=-2.5,hms_theta=35.0, |
||
| 1532 | 08:45 target changed to C12 |
||
| 1533 | target=Carbon,shms_p=-9.2,shms_theta=8,hms_p=-2.5,hms_theta=35.0, |
||
| 1534 | |||
| 1535 | 10:40 Beam is off |
||
| 1536 | |||
| 1537 | 3358 cosmic, 13:30 Oct 4th - 13:40 Oct 5th, no beam for the whole run |
||
| 1538 | |||
| 1539 | 3359 cosmic, 13:42 Oct 5th - 13:49 Oct 6th, no beam for the whole run |
||
| 1540 | |||
| 1541 | 3360 cosmic, 13:51 Oct 6th - 15:24 Oct 6th, no beam for the whole run |
||
| 1542 | |||
| 1543 | ==change TS#3 prescale factor to 1== |
||
| 1544 | |||
| 1545 | 3361 cosmic, 15:26 Oct 6th - 17:45 Oct 6th, no beam for the whole run |
||
| 1546 | |||
| 1547 | == enable TS#4, which is VME DSC ELC output ch4 == |
||
| 1548 | |||
| 1549 | 6 | Jixie Zhang | 3362 production, 17:47 Oct 6th - 01:02 Oct 10th, no beam until 03:22 Oct 9th |
| 1550 | 03:28 Oct 9th, target changed to Aluminum Hole,shms_p=-9.2,shms_theta=8,hms_p=-1.95,hms_theta=25.99, |
||
| 1551 | 04:16 Oct 9th, Target changed to LD2, |
||
| 1552 | 04:45 Oct 9th, Target changed to carbon, |
||
| 1553 | 05:53 Oct 9th, Target changed to LD2, |
||
| 1554 | 07:46 Oct 9th, Target changed to Carbon, |
||
| 1555 | 09:16 Oct 9th, Target changed to LD2, |
||
| 1556 | 10:26 Oct 9th, Target changed to Carbon, |
||
| 1557 | 11:47 Oct 9th, SHMS rotated to 8.5deg |
||
| 1558 | 13:40 Oct 9th, Target changed to LD2,shms_p=-9.2,shms_theta=8.50,hms_p=-1.95,hms_theta=25.99, |
||
| 1559 | 15:18 Oct 9th, SHMS rotated to 26deg,shms_p=-1.95,shms_theta=26.00,hms_p=-1.95,hms_theta=25.99 |
||
| 1560 | 15:55 Oct 9th, Target changed to LH2, |
||
| 1561 | 16:16 Oct 9th, Target changed to dummy, |
||
| 1562 | 17:36 Oct 9th, Target changed to Ca40, |
||
| 1563 | 19:00 Oct 9th, Target changed to Carbon, |
||
| 1564 | 5 | Jixie Zhang | beam stopped at 19:05 Oct 9th |
| 1565 | 1 | Jixie Zhang | |
| 1566 | 3363 production, 01:02 Oct 10th - 18:23 Oct 10th |
||
| 1567 | 6 | Jixie Zhang | target=Carbon,,shms_p=-1.95,shms_theta=26.00,hms_p=-1.95,hms_theta=25.99 |
| 1568 | 02:00 Oct 10th, Target changed to Dummy, |
||
| 1569 | 02:24 Oct 10th, Target changed to LD2, |
||
| 1570 | 02:51 Oct 10th, Target changed to LH2, |
||
| 1571 | 03:15 Oct 10th, both SHMS and HMS momenta changed to -5.42: |
||
| 1572 | target=LH2,shms_p=-5.42,shms_theta=26.00,hms_p=-5.42,hms_theta=25.99 |
||
| 1573 | 08:25 Oct 10th, Target changed to LD2, |
||
| 1574 | 17:18 Oct 10th, Target changed to Dummy, |
||
| 1575 | 1 | Jixie Zhang | |
| 1576 | 3364 production, 18:24 Oct 10th - 14:52 Oct 11th |
||
| 1577 | 6 | Jixie Zhang | target=Dummy,shms_p=-5.42,shms_theta=26.00,hms_p=-5.42,hms_theta=25.99, daq rate 6.4kHz @60uA |
| 1578 | 19:25 Oct 10th, Target changed to Ca40, |
||
| 1579 | 21:46 Oct 10th, Target changed to Carbon, |
||
| 1580 | 00:45 both SHMS and HMS momenta changed to -4.767 |
||
| 1581 | target=Carbon,shms_p=-4.767,shms_theta=26.00,hms_p=-4.767,hms_theta=25.99 |
||
| 1582 | 05:09 Oct 11th, Target changed to Ca40, |
||
| 1583 | 09:18 Oct 11th, Target changed to Dummy, |
||
| 1584 | 12:33 Oct 11th, Target changed to LD2, |
||
| 1585 | 13:49 Oct 11th, Target changed to LH2, |
||
| 1586 | 1 | Jixie Zhang | daq crashed at 06:07 Oct 11th |
| 1587 | |||
| 1588 | 6 | Jixie Zhang | 3365 production, 14:54 Oct 11th - 20:52 Oct 11th |
| 1589 | target=LH2,shms_p=-4.767,shms_theta=26.00,hms_p=-4.767,hms_theta=25.99 |
||
| 1590 | 16:45 both SHMS and HMS momenta changed to -4.19 |
||
| 1591 | target=LH2,shms_p=-4.19,shms_theta=26.00,hms_p=-4.19,hms_theta=25.99, |
||
| 1592 | 19:33 Oct 11th, Target changed to LD2, |
||
| 1593 | 1 | Jixie Zhang | |
| 1594 | 6 | Jixie Zhang | ==disable TS#4 @ 20221011 20:53== |
| 1595 | 3366 production, 20:58 Oct 11th - 13:28 Oct 13th |
||
| 1596 | target=LH2,shms_p=-4.19,shms_theta=26.00,hms_p=-4.19,hms_theta=25.99, daq rate 6.7kHz@60uA |
||
| 1597 | 21:12 Oct 11th, Target changed to Dummy, |
||
| 1598 | 23:10 Oct 11th, Target changed to Ca40, |
||
| 1599 | daq crashed at 21:32 Oct 11th |
||
| 1600 | 1 | Jixie Zhang | |
| 1601 | 6 | Jixie Zhang | 3367 pedestal, 13:32 Oct 13th - 13:44 Oct 13th, |
| 1602 | target=Carbon, shms_p=2.52,shms_theta=26.00,hms_p=2.52,hms_theta=25.99, |
||
| 1603 | 13:38 target changed to dummy |
||
| 1604 | daq crashed |
||
| 1605 | |||
| 1606 | 3368 production, 13:45 Oct 13th - 14:22 Oct 13th |
||
| 1607 | target=Dummy, shms_p=2.52,shms_theta=26.00,hms_p=2.52,hms_theta=25.99, |
||
| 1608 | daq crashed |
||
| 1609 | |||
| 1610 | 3369 production, 14:25 Oct 13th - 14:45 Oct 13th |
||
| 1611 | target=Dummy, shms_p=2.52,shms_theta=26.00,hms_p=2.52,hms_theta=25.99, daq rate 6.2kHz@58uA |
||
| 1612 | daq crashed |
||
| 1613 | 1 | Jixie Zhang | |
| 1614 | 6 | Jixie Zhang | 3370 production, 14:52 Oct 13th - 17:55 Oct 13th |
| 1615 | 14:53 target changed to LD2 |
||
| 1616 | target=LD2, shms_p=2.52,shms_theta=26.00,hms_p=2.52,hms_theta=25.99, daq rate 6.6kHz@58uA |
||
| 1617 | 15:24 target changed to LH2 |
||
| 1618 | 16:12 target changed to Ca40 |
||
| 1619 | 16:15 both SHMS and HMS momenta changed to -2.21 |
||
| 1620 | target=Ca40, shms_p=2.21,shms_theta=26.00,hms_p=2.21,hms_theta=25.99, |
||
| 1621 | 17:05 target changed to Carbon |
||
| 1622 | |||
| 1623 | 3371 production, 17:55 Oct 13th - 23:59 Oct 13th |
||
| 1624 | target=Carbon, shms_p=2.21,shms_theta=26.00,hms_p=2.21,hms_theta=25.99, daq rate 6.7kHz@60uA |
||
| 1625 | |||
| 1626 | 3372 production, 23:59 Oct 13th - 15:02 Oct 14th |
||
| 1627 | |||
| 1628 | beam stopped at 8 AM |
||
| 1629 | |||
| 1630 | |||
| 1631 | 3373 production, 15:03 Oct 14th - 02:16 Oct 15th |
||
| 1632 | daq crashed at 18:35 Oct 14th due to disk is full |
||
| 1633 | |||
| 1634 | |||
| 1635 | 3374 trash, 02:18 Oct 15th - daq could not start |
||
| 1636 | 3375 trash, 02:22 Oct 15th - daq could not start |
||
| 1637 | |||
| 1638 | 3376 production, 02:28 Oct 15th - 16:48 Oct 15th |
||
| 1639 | daq rate 6.4kHz @ 40uA |
||
| 1640 | daq crashed at 14:14 Oct 15th due to disk is full |
||
| 1641 | |||
| 1642 | 3377 trash, 17:20 Oct 15th - daq could not start |
||
| 1643 | |||
| 1644 | 3378 production, 17:21 Oct 15th - 13:55 Oct 16th |
||
| 1645 | daq rate 6.6kHz @ 59uA |
||
| 1646 | daq crashed sincd 05:42, no beam since 08:55 Oct 16th |
||
| 1647 | |||
| 1648 | 3379 production, 17:58 Oct 16th - 15:26 Oct 17th |
||
| 1649 | daq crashed sincd 12:47, |
||
| 1650 | |||
| 1651 | 3380 production, 15:29 Oct 17th - 00:45 Oct 18th |
||
| 1652 | daq crashed @ 22:45 |
||
| 1653 | |||
| 1654 | 3381 production, 00:47 Oct 18th - 00:34 Oct 19th |
||
| 1655 | daq rate 4.94kHz @59 uA |
||
| 1656 | no beam since 03:27 |
||
| 1657 | |||
| 1658 | 3382 production, 00:34 Oct 19th - 13:37 Oct 19th |
||
| 1659 | no beam since 08:00 |
||
| 1660 | |||
| 1661 | 3383 production, 23:54 Oct 19th - 10:59 Oct 20th |
||
| 1662 | 10:58 daq rate 1.250Hz @ 45uA |
||
| 1663 | |||
| 1664 | 3384 production, 11:00 Oct 20th - ? Oct 20th |
||
| 1665 | 11:02 daq rate 1.250Hz @ 45uA |